### v3

# Dynamical Moiré Systems in Twisted Bilayer Optical Lattices

Abstract
Moiré related physics in twisted bilayer two-dimensional (2D) materials has attracted widespread interest in condensed matter physics. Simulation of moiré related physics in cold atom platform is expected to outperform the 2D materials thanks to its advantage of higher tunablility. Here, we point out, the cold atom platform enables a new mechanism of moiré lattice formation, with intrinsic “dynamical” character arising from interlayer interaction, in contrast to conventional moiré lattice induced by “static” ways such as single-particle interlayer tunneling or external controls. Specifically, we consider a twisted bilayer Bose-Hubbard model with vanishing interlayer tunneling, and the bilayer is solely coupled through interlayer interaction that originates from contact interaction of atoms. We find that this system hosts a plethora of novel phases unique to this dynamical lattice, including a variety of Mott insulator (MI) and superfluid (SF) phases either preserving or breaking moiré lattice symmetry, phases with one layer in SF and the other in MI, “interlocked” MI, and self-localized phases at commensurate twist angles, which exhibits the characteristics of Bose glass and quasi-many-body localization in the absence of (quasi)disorder or quasicrystalline lattices. Our prediction can be readily observed in current experimental setup of twisted bilayer optical lattices, opening up new avenues for exploring the rich physics of dynamical moiré systems in cold atoms.
Introduction.—Moiré physics has been extensively studied in the field of condensed matter physics in recent years. One prominent feature that stimulates tremendous interest is the emergence of flat band by simply tuning the twist angle of bilayer two-dimensional (2D) materials, which offers an intriguing platform for exploring phases that electron-electron interaction plays a significant role, such as correlated insulators Cao et al. (2018a); Li et al. (2021a), superconductivity Cao et al. (2018b), magnetism Gong et al. (2017); Huang et al. (2017); Chen et al. (2020) and generalized Wigner crystals Regan et al. (2020); Li et al. (2021b). On the other hand, topological phases also emerge in moiré superlattices, e.g., quantum anomalous Hall insulators (QAHI) Serlin et al. (2020); Li et al. (2021c); Zhou et al. (2021), $\mathbb{Z}_{2}$ topological insulators Wu et al. (2019); Zhao et al. (2024), high-order topological insulators Park et al. (2019), and fractional QAHI Cai et al. (2023); Park et al. (2023); Zeng et al. (2023); Xu et al. (2023); Lu et al. (2024). The rich many-body and topological physics observed in bilayer 2D materials have clearly demonstrated that twisting serves as a novel and effective means for manipulating material properties and unlocking diverse and fascinating quantum phenomena Kennes et al. (2021).
Exploring the moiré physics in other platforms is appealing to fully unleash the power of twisting. One prospective platform is cold atoms, where optical lattice has been a faithful simulation of lattice physics in condensed matter Bloch et al. (2008). A variety of theoretical proposals have been put forward González-Tudela and Cirac (2019); Salamon et al. (2020); Luo and Zhang (2021); Paul et al. (2023); Wan et al. (2024); Wang et al. (2024); Madroñero and Paredes (2023); Madroñero et al. (2024), where twisted moiré physics can be either simulated using bilayer optical lattice or using pseudospin-1/2 atomic species. Very recently, exciting experimental progress has been made Meng et al. (2023), which has turned long-standing speculation into reality. Compared to 2D materials, cold atom platforms offer significant advantages Bloch et al. (2008); Chin et al. (2010); Gross and Bakr (2021), as they allow for the engineering of diverse monolayer lattice structures, dynamic tuning of twist angles, and straightforward adjustment of interlayer coupling. Moreover, being free from lattice relaxation, strain, and disorder, these platforms support pristine twisted bilayer models, allowing direct comparisons between theories and experiments.
So far, the moiré lattices observed in both 2D materials and cold atoms share a common feature: they are “static”, in the sense that their forms are uniquely fixed by deterministic factors such as interlayer tunneling, structural relaxations, or external controls like dielectric screening Bistritzer and MacDonald (2011); Wang et al. (2017); Ruiz-Tijerina and Fal’ko (2019); Li et al. (2021d); Xu et al. (2021); Kim et al. (2024); Zhang et al. (2024); Gu et al. (2024); He et al. (2024). In this work, we propose a distinct mechanism of realizing moiré lattice solely induced by particle-particle interaction from two layers or pseudospins, which we refer to as “dynamical” moiré lattice  [^1] This is to be contrasted with dynamical control of moiré lattice already realized in 2D materials, where the moiré potential in the target layer is uniquely imprinted by substrate layers and has no feedback on the substrate layers Kim et al. (2024); Zhang et al. (2024); Gu et al. (2024); He et al. (2024). For each layer/pseudospin of lattice, the effect of the other layer is brought by particle-particle interaction, which relies on particle density distribution, an intrinsic dynamical variable. Therefore, each layer can feel dynamical feedback from the other layer and the ultimate density distribution is a compromise between the two layers. Realizing such a dynamical moiré system in twisted 2D materials is challenging, but it can be easily achieved in current cold atom platforms by simply turning off interspin coupling in twisted spin-dependent lattices Meng et al. (2023).
Figure 1: (a) Schematic illustration of the dynamical moiré system. (b) Sketch of a twisted bilayer optical lattice at the twisted angle $\theta=2\arctan(1/3)=36.87^{\circ}$. Red (blue) triangles correspond to lattice sites of $\uparrow$ ($\downarrow$) layer. The purple circle denotes the external cylinder box trap. (c) Schematic illustration of the origin of the interlayer interaction. Filled triangles denote the lattice sites occupied by atoms. The overlap between atomic wave functions (denoted by the translucent circles) of atoms on different layers gives rise to the interlayer interaction. (d) Zoo of rich phases in the dynamical moiré system. The left (right) column of each subplot corresponds to the ($\uparrow$($\downarrow$)↑ ( ↓ )-layer, respectively. From left to right, up to down, ($\theta$, $\rho_{\uparrow}$, $\rho_{\downarrow}$, $g_{\uparrow\downarrow}/g_{\sigma\sigma}$, $J/U(g_{\sigma\sigma})$) assume the values ($7.63^{\circ}$, $3/5$, $3/5$, $0.5$, $0.06$), ($36.87^{\circ}$, $4/5$, $4/5$, $0.9$, $0.048$), ($36.87^{\circ}$, $7/10$, $4/5$, $0.9$, $0.04$), ($22.62^{\circ}$, $2/5$, $3/5$, $0.9$, $0.024$), ($36.87^{\circ}$, $4/5$, $4/5$, $0.3$, $0.008$), and ($7.63^{\circ}$, $7/10$, $3/5$, $0.5$, $0.012$), respectively. See text for more details.
Specifically, we consider two species of ultracold atoms loaded in a twisted bilayer spin-dependent optical lattice and find that, an emergent moiré lattice from particle-interaction can form, hosting much more rich phases than static lattice counterpart (see Fig. 1), including phases with and without moiré lattice symmetry and Mott insulator (MI) phases induced by interlayer interaction even for partial filling. More interestingly, for total unit filling, the ground state can become an “interlocked” MI with spontaneously formed moiré lattice. For a general total filling, the Bose glass phase exists, suggesting that the ground state tends to localize due to the interplay between twisting and interlayer interaction, even at commensurate twist angles.
System and model.—Motivated by the recent realization of the twisted bilayer optical lattices Meng et al. (2023), we consider a closely related but simpler experimental set-up, namely, a twisted bilayer optical lattice system without the additional microwave field for realizing the interlayer coupling in the recent experiments Meng et al. (2023). More specifically, the system consists of two species of bosonic atoms that correspond to two hyperfine states (for instance, ${}^{87}\mathrm{Rb}$ atoms in hyperfine states $|F=1,m_{F}=1\rangle$ and $|F=2,m_{F}=0\rangle$ considered in Ref. Meng et al. (2023)), with each species loaded in one layer of a spin-dependent twisted bilayer square optical lattice, respectively [see Fig. 1(a)].
The physics of this system at low filling can be described by a Bose-Hubbard type model within the lowest band approximation (see Supplemental Material (SM) Sup for a detailed derivation),
$\hat{H}=$ $\sum_{\sigma=\uparrow,\downarrow}\left[-J\sum_{\langle\mathbf{i}_{\sigma},\mathbf{j}_{\sigma}\rangle}\hat{b}_{\mathbf{i_{\sigma}}}^{\dagger}\hat{b}_{\mathbf{j_{\sigma}}}+\sum_{\mathbf{i}_{\sigma}}\frac{U(g_{\sigma\sigma})}{2}\hat{n}_{\mathbf{i}_{\sigma}}(\hat{n}_{\mathbf{i}_{\sigma}}-1)\right]$
$+\sum_{\sigma=\uparrow,\downarrow}\sum_{\langle\mathbf{i}_{\sigma},\mathbf{j}_{\sigma}\rangle}\frac{U_{\mathbf{i}_{\sigma}\mathbf{j}_{\sigma}}(g_{\sigma\sigma})}{2}\hat{n}_{\mathbf{i}_{\sigma}}\hat{n}_{\mathbf{j}_{\sigma}}$
$+\sum_{\langle\mathbf{i}_{\uparrow},\mathbf{j}_{\downarrow}\rangle}U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}(g_{\uparrow\downarrow},\theta)\hat{n}_{\mathbf{i}_{\uparrow}}\hat{n}_{\mathbf{j}_{\downarrow}},$ (1)
[…]
The first two parts of $\hat{H}$ basically assume the form of the conventional Bose-Hubbard model, which consists of the hopping term of different species of atoms in its own layer with the hopping amplitude being $J$ and the interaction terms originating from the intra-species contact interactions with strength $g_{\sigma\sigma}$ between atoms on the same layer. For the latter, we not only take into account the conventional on-site interaction term with the strength denoted by $U(g_{\sigma\sigma})$, but also the nearest-neighbor interaction terms with the strengths denoted by $U_{\mathbf{i}_{\sigma}\mathbf{j}_{\sigma}}(g_{\sigma\sigma})$, and treat the inter-species contact interaction on an equal footing.
The third part of $\hat{H}$ captures the distinct way that the twist imposes its physical influences even without the interlayer tunneling. It originates from the finite overlaps between wave functions of atoms on different layers with their inter-species contact interaction being $g_{\uparrow\downarrow}$ [see Fig. 1(c)], and consists of interlayer interaction terms $U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}(g_{\uparrow\downarrow},\theta)\hat{n}_{\mathbf{i}_{\uparrow}}\hat{n}_{\mathbf{j}_{\downarrow}}$, with the interaction strength $U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}(g_{\uparrow\downarrow},\theta)$ assuming the twist angle $\theta$ dependence. This dependence reflects how the overlaps between Wannier wave functions of different sites on different layers are changed when one tunes $\theta$.
Figure 2: Phase diagrams and typical configurations at the balanced and imbalanced fillings. (a) Phase diagram with respect to the interlayer interaction strength $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ and the hopping amplitude $J/U(g_{\sigma\sigma})$ at the balanced filling with $\rho_{\uparrow}=4/5$ and $\rho_{\downarrow}=4/5$ (twist angle is fixed at $\theta=36.87^{\circ}$). (b) Similar to (a), but at the imbalanced filling with $\rho_{\uparrow}=1/5$ and $\rho_{\downarrow}=4/5$. (c) and (d) Real-space particle density and superfluid order parameter distributions of the system that correspond to the red dots and the blue squares in (a) and (b), respectively. See text for more details.
We investigate this system by calculating its ground state via the bosonic Gutzwiller variational approach Krauth et al. (1992); Jaksch et al. (1998); Lanatà et al. (2012) (see SM Sup for technical details). We have chosen the twist angle $\theta=2\arctan(\bar{m}/\bar{n})$ with $\bar{m},\bar{n}$ being two natural numbers Wang et al. (2020); Meng et al. (2023), i.e., the commensurate condition for twisted square lattice, and focus on the case with $g_{\uparrow\uparrow}=g_{\downarrow\downarrow}$. If not specified otherwise, we perform the numerical simulation on a $37\times 37$ twisted bilayer optical lattice with an external 2D cylinder box trap Navon et al. (2021) whose radius is 18 lattice constants (denoted by the solid circles in the figures). In the following, we focus on the most interesting case of fractional filling for both layers, and reserve the results of integer filling in the Supplemental Material Sup (the impact of interlayer interaction is not as dramatic in the latter case).
Before presenting the numerical results, we first notice that in conventional moiré systems Bistritzer and MacDonald (2011); Wang et al. (2017); Ruiz-Tijerina and Fal’ko (2019); Meng et al. (2023), one layer of lattice actually influences the physics on the other layer in a “static” way, in the sense that the influences are solely determined by the configuration of the twisted bilayer lattice through the single-particle interlayer tunneling. In sharp contrast, here, these influences are “dynamical”, in the sense that one layer of lattice impose its influences on the physics of the other via the dynamical variables on this layer (particle number operators in this case), as clearly manifested by the interlayer interaction terms of the Hamiltonian (1), giving rise to a new type of moiré system we dubbed as “dynamical moiré system”. This intrinsic dynamical character could make the system feature not only similar phases in the conventional moiré system, but also many intriguing phases without counterpart in the latter [see Fig. 1(d) for an overview].
Conventional moiré physics without interlayer tunneling.—In conventional bosonic moiré systems, the interlayer tunneling causes layer hybridization of single particle orbitals, and thus particle filling on the hybridized orbitals results in phases with two layer sharing the same character. Indeed, as observed in recent experiments Meng et al. (2023), it can accommodate SF-SF and MI-MI (denoting SF/MI phase in $\uparrow$/$\downarrow$ / ↓ layer), with the particle density assuming the moiré lattice symmetry. In fact, these observations can be easily explained by treating the system as if the atoms on one layer experience an additional static potential from the other layer of the lattice Meng et al. (2023). Speculating along this line, we notice that although there is no interlayer tunneling in the dynamical moiré system studied here, atoms residing on one layer can nevertheless feel an additional potential provided by the atoms residing on the other layer via the interlayer interactions $\sum_{\langle\mathbf{i}_{\uparrow},\mathbf{j}_{\downarrow}\rangle}U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}(g_{\uparrow\downarrow},\theta)\hat{n}_{\mathbf{i}_{\uparrow}}\hat{n}_{\mathbf{j}_{\downarrow}}$. This indicates the system may accommodate similar conventional moiré physics even without the interlayer tunneling.
In Fig. 2(a), we map out the phase diagram of the system in a balanced filling case with $\rho_{\uparrow}=\rho_{\downarrow}=4/5$ ($\rho_{\sigma}\equiv\langle\sum_{\mathbf{i}_{\sigma}}\hat{n}_{\mathbf{i}_{\sigma}}\rangle/N_{\sigma}^{\mathrm{lat}}$ with $N_{\sigma}^{\mathrm{lat}}$ being the number of lattice sites of the $\sigma$-layer within the cylinder trap). In the relatively large hopping regime, we can see that the system indeed support SF-SF phase (denoted as SF in the phase diagrams) similar to conventional static moiré systems, with both layers manifesting the SF order parameter $\phi_{\mathbf{i}_{\sigma}}\equiv\langle\hat{b}_{\mathbf{i}_{\sigma}}\rangle$ and density distribution $\langle\hat{n}_{\mathbf{i}_{\sigma}}\rangle$ assuming the moiré lattice symmetry [see the first plot in Fig. 2(c)]. While in the relatively small hopping regime, we notice that despite the filling factor on each layer being non-integer, increasing the inter-species interaction strength $g_{\uparrow\downarrow}$ can drive the system into the induced MI phase characterized by regions with integer-filled particles and vanishing SF order parameter in either layer [see the last plot in Fig. 2(c)], including MI-SF, MI-MI and SF-MI phases. This is in sharp contrast to the single-component Bose gases in optical lattices, where the MI can only exist at integer fillings.
The induced MI phase occurs because increased $g_{\uparrow\downarrow}$ enhances the effective potential experienced by atoms on one layer imposed by atoms on the other layer, causing localization. This scenario is reminiscent of conventional bosonic moiré systems, where increasing the interlayer tunneling makes one layer impose a stronger additional static lattice potential on the other layer, leading to stronger localization Meng et al. (2023). However, it’s important to recognize a key difference here: the effective potential is a dynamical one determined by the dynamical variables, and hence does not necessarily share the same spatial symmetry of the underlying optical lattice. For balanced filling case, indeed, we have not found MI phase with both layer assumes the moiré lattice symmetry. Instead, as shown in the second and third plot in Fig. 2(c), we find this dynamical moiré system can manifest rich discrete rotational symmetric MI and SF phases with no counterparts in the conventional bosonic moiré systems. However, away from the balanced filling, we indeed find the system can support MI phase with density distributions on both layers assuming the moiré lattice symmetry. For instance, as shown in the last two plots in Fig. 2(d), this kind of moiré MI phase can emerge in the induced MI regime of the system at an imbalanced filling $\rho_{\uparrow}=1/5,\rho_{\downarrow}=4/5$.
Although the dynamical moiré system can manifest SF or MI phase similar to the ones found in conventional bosonic moiré systems Meng et al. (2023), there are fundamental differences. Instead of being coupled by interlayer tunneling, the atoms on different layers are coupled by the interlayer density-density interactions. Consequentially, the total particle number of each layer is conserved and can be different. These fundamental differences thus naturally indicate much richer physics with no counterparts in conventional bosonic moiré systems can emerge, as we shall discuss below.
Beyond the conventional moiré physics.—By comparing the physics at balanced and imbalanced filling as showcased in Fig. 2, one clearly notices the strong physical influences of the filling difference. Since the total particle number of each layer is conserved separately, the filling difference between the two layers in fact provides an extra tuning knob of new physics compared with the conventional bosonic moiré systems. In Fig. 3(a), we systematically investigate the influences of the filling difference by keeping the filling factor of the $\downarrow$-layer fixed at $\rho_{\downarrow}=4/5$ and changing the filling factor of the $\uparrow$-layer $\rho_{\uparrow}$ from $1/10$ all the way to $4/5$. As we can see from the first plot in Fig. 3(c), when the filling of the $\uparrow$-layer is small ($\rho_{\uparrow}=1/10$), the $\downarrow$-layer is in a SF phase and it imposes strong influences on atoms on the $\uparrow$-layer by localizing them and making them in induced MI phase despite $\rho_{\uparrow}=1/10$ being non-integer. As the $\rho_{\uparrow}$ gradually increases [see the last two plots in Fig. 3(c)], the $\uparrow$-layer also imposes an enhanced localization effects on the atoms on the $\downarrow$-layer by driving a SF-MI transition on this layer. We also observe that the average localization effect experienced by each atom in the $\uparrow$-layer decreases, making the $\uparrow$-layer transition into a SF phase at relatively high filling with $\rho_{\uparrow}=1/2$. This phenomenon is reminiscent of what happens in high-temperature superconductors, where doping can lead to significant changes in the system’s transport properties Keimer et al. (2015). More interestingly, for the total unit filling, the system hosts a kind of “interlocked” MI with “interlocked” density distributions [see the second plot of Fig. 3(c), the lower left plot of Fig. 1(d) and the fourth plot of Fig. 2(d)].
Similar to the conventional moiré system, the twist angle also has a significant impact here, as clearly shown in Fig. 3(b). For the case at the filling $(\rho_{\uparrow}=1/10,\rho_{\downarrow}=3/5)$, in the intermediate hopping regime, we see that at a relatively small twist angle $\theta=7.63^{\circ}$, the system is in the SF-SF phase. As the twist angle increases to $\theta=22.62^{\circ}$, the atoms in the $\uparrow$-layer are localized due to the interaction from the $\downarrow$-layer and enter the induced MI phase, forming the phase MI-SF [see the second plot of Fig. 3(d)]. Similar situation is also found at even larger twist angle, for instance, $\theta=36.87^{\circ}$ [see the last plot of Fig. 3(d)].
Interestingly, we notice that density distributions in the last two plots of Fig. 3(d) assume rather random pattern. In these cases, the atoms on $\downarrow$-layer in fact provide a “disordered” effective potential to the atoms lying on the $\uparrow$-layer. Thus, one naturally expects that this “disordered” effective potential quenches hopping of the $\uparrow$-atoms and makes them localized, which can be interpreted as the interlayer interaction induced “self-localization” between the atoms on the two layers. Among the states with irregular density pattern, such as the third plot of Fig. 3(d) and the last plot of Fig. 1(d), there exist states with disconnected superfluid islands, exhibiting the characteristics of Bose glass. Thus, this system supports Bose glass without extrinsic disorder, in contrast to previous studies with either (quasi)disorder or quasicrystalline lattices Fisher et al. (1989); Fallani et al. (2007); Pasienski et al. (2010); Söyler et al. (2011); Gadway et al. (2011); D’Errico et al. (2014); Meldgin et al. (2016); Johnstone et al. (2021); Ciardi et al. (2023). The presence of Bose glass is associated with the spontaneous breaking of $C_{4}$ rotational symmetry of the system, indicating that there are four degenerate ground states. As the presence of Bose glass at ground state usually implies many-body localization (MBL) at excited states Abanin et al. (2019), this system also probably hosts (quasi-)MBL without extrinsic disorder Schiulaz et al. (2015); Yao et al. (2016); Smith et al. (2017); Karpov et al. (2021). Note that here either the Bose glass or the possible quasi-MBL are found at commensurate twist angles, i.e., the system has translational invariance in the bulk, in contrast to previously investigated localization at incommensurate angles Huang and Liu (2019); Paul et al. (2023); Madroñero et al. (2024) or the interacting 2D Aubry-André model with quasi-periodicity Szabó and Schneider (2020); Štrkalj et al. (2022). The possible quasi-MBL here may be explained by a mechanism similar to the exponentially slow relaxation predicted in certain translation-invariant systems Schiulaz et al. (2015); Yao et al. (2016). Further exploration of the relation between the localization found here with MBL and the generalization to incommensurate twist angles will be an intriguing subject for future study.
Figure 3: (a) Phase diagram with respect to $\rho_{\uparrow}$ and $J/U(g_{\sigma\sigma})$ with other parameters fixed at $g_{\uparrow\downarrow}/g_{\sigma\sigma}=0.9$, $\theta=36.87^{\circ}$, and $\rho_{\downarrow}=4/5$. (b) Phase diagram with respect to the twist angle $\theta$ and $J/U(g_{\sigma\sigma})$ with other parameters fixed at $g_{\uparrow\downarrow}/g_{\sigma\sigma}=0.9$, $\rho_{\uparrow}=1/10$, and $\rho_{\downarrow}=3/5$. (c) and (d) Real-space particle density and superfluid order parameter distributions of the system that correspond to the red dots and the blue squares in (a) and (b), respectively. See text for more details.
In light of the tremendous experimental progress in ultracold atomic physics, the rich physics in this dynamical moiré system can be readily observed using current experimental setups Meng et al. (2023). Specifically, $J/U(g_{\sigma\sigma})$ can be tuned by adjusting the depth of optical lattice potential. The interlayer interaction $U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}(g_{\uparrow\downarrow},\theta)$ can be altered by either tuning the vertical displacement between the two spin-dependent lattices, or by utilizing the Feshbach resonance  [^2] The Feshbach resonance technique enables the direct tuning of $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ in the phase diagrams. For the interlayer interaction $U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}\equiv\frac{g_{\uparrow\downarrow}}{2}\int d^{3}\mathbf{r}\left|w_{\mathbf{i}_{\uparrow}}(\mathbf{r})\right|^{2}\left|w_{\mathbf{j}_{\downarrow}}(\mathbf{r})\right|^{2}$, displacing the spin-dependent lattices in out-of-plane direction reduces the overlap between Wannier functions $w_{{\uparrow/\downarrow}}(\mathbf{r})$, whose effect is physically equivalent to the change of $g_{\uparrow\downarrow}/g_{\sigma\sigma}$. The control over the twist angle can be realized by experimental techniques developed recently for the twisted bilayer optical lattices Meng et al. (2023). The exotic SF, MI phases, and various self-localization behavior can be directly observed via the quantum gas microscope Gross and Bakr (2021).
Conclusions.—We propose a new type of moiré system in twisted bilayer optical lattices, namely, the dynamical moiré system that hosts a wide range of unique physical phenomena. It can not only accommodate moiré SF phase similar to the one found in conventional static bosonic moiré system, but also give rise to rich physics without static counterparts. These include “interlocked” MI, exotic SF and MI phases breaking moiré lattice symmetry, phases with one layer in SF and the other in MI, localization behavior similar to Bose glass and quasi-MBL without extrinsic disorder. Since both the dynamical moiré system itself and its experimental realization is even simpler than the conventional bosonic moiré system realized in experiments recently Meng et al. (2023), we believe our work will stimulate both further theoretical research on this type of dynamical moiré system, especially the possible intrinsic connection between the self-localization found here and quasi-MBL, generalization to fermionic systems, as well as experimental effort in searching for these exotic quantum phases.
Acknowledgements.
This work is supported by NKRDPC (Grant No. 2022YFA1405304), NSFC (Grant Nos. 12004118 and 12275089), Guangdong Basic and Applied Basic Research Foundation (Grants Nos. 2020A1515110228, 2021A1515010212 and 2023A1515012800), Guangdong Provincial Key Laboratory (Grant No. 2020B1212060066), and START grant of South China Normal University.
Supplemental Material for “Dynamical Moiré Systems in Twisted Bilayer Optical Lattices”
I System and model
Motivated by the recent realization of the twisted bilayer optical lattices (Meng et al., 2023), we consider a closely related but simpler experimental set-up, namely, a twisted bilayer optical lattice system without the additional microwave field that has been used to realize the interlayer coupling in the recent experiments (Meng et al., 2023). More specifically, the system under study consists of two species of bosonic atoms that correspond to two hyperfine states (for instance, ${}^{87}\mathrm{Rb}$ atoms in hyperfine states $|F=1,m_{F}=1\rangle$ and $|F=2,m_{F}=0\rangle$ considered in Ref. (Meng et al., 2023)), with each species loaded in one layer of a state-dependent twisted bilayer square optical lattice, respectively [see Fig. 1(a) of the main text for a schematic illustration].
The physics of this system at low filling can be described by a Bose-Hubbard type model within the lowest band approximation. To derive the model Hamiltonian, we start with the following many-body Hamiltonian
$\hat{H}=$ $\sum_{\sigma=\uparrow,\downarrow}\int d^{3}\mathbf{r}\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r})\left(\frac{-\hbar^{2}}{2\mu}\nabla^{2}+V(\mathrm{\bm{r}})\right)\hat{\psi_{\sigma}}(\mathbf{r})$
$+\sum_{\sigma=\uparrow,\downarrow}\int d^{3}\mathbf{r}\int d^{3}\mathbf{r}^{\prime}\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r})\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r^{\prime}})\frac{V_{\sigma\sigma}(\mathbf{r}-\mathbf{r^{\prime}})}{2}\hat{\psi}_{\sigma}(\mathbf{r^{\prime}})\hat{\psi}_{\sigma}(\mathbf{r})$ (S1)
$+\int d^{3}\mathbf{r}\int d^{3}\bm{\mathbf{r}}^{\prime}\hat{\psi}_{\uparrow}^{\dagger}(\mathbf{r})\hat{\psi}_{\downarrow}^{\dagger}(\mathbf{r^{\prime}})V_{\uparrow\downarrow}(\mathbf{r}-\mathbf{r^{\prime}})\hat{\psi}_{\downarrow}(\mathbf{r^{\prime}})\hat{\psi}_{\uparrow}(\mathbf{r}),$
where $\sigma=\uparrow,\downarrow$ is the index of the atom species and lattice layer, $V_{\sigma\sigma}(\mathbf{r}-\mathbf{r^{\prime}})\left[V_{\uparrow\downarrow}(\mathbf{r}-\mathbf{r^{\prime}})\right]$ represents the interaction between two intra(inter)-species atoms, $V(\mathrm{\bm{r}})$ represents the lattice potential, $\mu$ is the mass of an atom, and the operators $\hat{\psi_{\sigma}}^{\dagger}(\mathbf{r})$ and $\hat{\psi_{\sigma}}(\mathbf{r})$ respectively create and annihilate an atom at position $\bm{\mathbf{r}}$ of the lattice layer $\sigma$. In the context of ultracold atomic gases, where the interactions are modeled by effective two-particle interactions for suciently dilute gases, a choice suitable for most theoretical models is the delta contact potential between two atoms of mass $\mu$ at positions $\mathbf{r}$ and $\mathbf{r^{\prime}}$
$V_{\sigma\sigma}(\mathbf{r}-\mathbf{r^{\prime}})$ $=\frac{4\pi\hbar^{2}a_{\sigma\sigma}}{\mu}\delta(\mathbf{r}-\mathbf{r^{\prime}}),$
$V_{\uparrow\downarrow}(\mathbf{r}-\mathbf{r^{\prime}})$ $=\frac{4\pi\hbar^{2}a_{\uparrow\downarrow}}{\mu}\delta(\mathbf{r}-\mathbf{r^{\prime}}),$ (S2)
with $a_{\sigma\sigma}$ ($a_{\uparrow\downarrow}$) being the intra(inter)-species $s$-wave scattering length. The first term of the Hamiltonian (S1) is the single particle term, and the second term describes the intra-species interaction of atoms in the same lattice layer, and the third term describes the inter-species interaction between atoms of different lattice layer.
Within the second quantization formalism, we can define the creation operator for the lowest band Wannier orbital $w_{\mathbf{i}_{\sigma}}(\mathbf{r})=\langle\mathbf{r}\mid\mathbf{i}_{\sigma}\rangle$ at site $\mathbf{i}$ of $\sigma$-layer,
$\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}=\int d^{3}\mathbf{r}w_{\mathbf{i}_{\sigma}}(\mathbf{r})\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r}).$ (S3)
Therefore,
$\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r})=\sum_{\mathbf{i}_{\sigma}}\langle\mathbf{i}_{\sigma}\mid\mathbf{r}\rangle\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}=\sum_{\mathbf{i}_{\sigma}}w_{\mathbf{i}_{\sigma}}^{*}(\mathbf{r})\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger},$ (S4)
$\hat{\psi}_{\sigma}(\mathbf{r})=\sum_{\mathbf{i}_{\sigma}}\langle\mathbf{r}\mid\mathbf{i}_{\sigma}\rangle\hat{b}_{\mathbf{i}_{\sigma}}=\sum_{\mathbf{i}_{\sigma}}w_{\mathbf{i}_{\sigma}}(\mathbf{r})\hat{b}_{\mathbf{i}_{\sigma}},$ (S5)
where $\hat{b}_{\mathbf{i_{\sigma}}}^{\dagger}(\hat{b}_{\mathbf{i}_{\sigma}})$ is the creation (annihilation) operator at site $\mathbf{i}$ of the lattice layer $\sigma$ in the Wannier basis.
For a relatively shallow optical lattice, the single particle Hamiltonian within the tight-binding approximation restricted to lowest band reads
$H_{\mathrm{s.p.},\sigma}=$ $\int d^{3}\mathbf{r}\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r})\left(\frac{-\hbar^{2}}{2\mu}\nabla^{2}+V(\mathrm{\bm{r}})\right)\hat{\psi_{\sigma}}(\mathbf{r})$
$=$ $\sum_{\mathbf{i}_{\sigma},\mathbf{j}_{\sigma}}\int d^{3}\mathbf{r}w_{\mathbf{i}_{\sigma}}^{*}(\mathbf{r})\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\left(\frac{-\hbar^{2}}{2\mu}\nabla^{2}+V(\mathrm{\bm{r}})\right)w_{\mathbf{j}_{\sigma}}(\mathbf{r})\hat{b}_{\mathbf{j}_{\sigma}}$ (S6)
$=$ $\sum_{\mathbf{i}_{\sigma}}\epsilon_{\mathbf{i}_{\sigma}}\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{i}_{\sigma}}-J\sum_{\left\langle\mathbf{i}_{\sigma},\mathbf{j}_{\sigma}\right\rangle}\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{j}_{\sigma}},$
[…]
$\epsilon_{\mathbf{i}_{\sigma}}\equiv\int d^{3}\mathbf{r}w_{\mathbf{i}_{\sigma}}^{*}(\mathbf{r})[\frac{-\hbar^{2}}{2\mu}\nabla^{2}+V(\mathrm{\bm{r}})]w_{\mathbf{i}_{\sigma}}(\mathbf{r}),$ (S7)
and
$J\equiv-\int d^{3}\mathbf{r}w_{\mathbf{i}_{\sigma}}^{*}(\mathbf{r})[\frac{-\hbar^{2}}{2\mu}\nabla^{2}+V(\mathrm{\bm{r}})]w_{\mathbf{j}_{\sigma}}(\mathbf{r}).$ (S8)
For a sufficiently dilute, weakly interacting atomic gas, the short ranged interactions covering the nearest neighbor sites of single component are taken into consideration. The interaction Hamiltonian of single component reads
$H_{\mathrm{intra},\sigma}=$ $\int d^{3}\mathbf{r}\int d^{3}\mathbf{r}^{\prime}(\mathbf{r})\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r})\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r^{\prime}})\frac{V_{\sigma\sigma}(\mathbf{r}-\mathbf{r^{\prime}})}{2}\hat{\psi}_{\sigma}(\mathbf{r^{\prime}})\hat{\psi}_{\sigma}(\mathbf{r})$
$=$ $\sum_{\mathbf{i}_{\sigma},\mathbf{j}_{\sigma},\mathbf{k}_{\sigma},\mathbf{l}_{\sigma}}\frac{1}{2}\int d^{3}\mathbf{r}\int d^{3}\mathbf{r}^{\prime}w_{\mathbf{i}_{\sigma}}^{*}(\mathbf{r})w_{\mathbf{j}_{\sigma}}^{*}(\mathbf{r^{\prime}})\frac{4\pi\hbar^{2}a_{\sigma\sigma}}{\mu}\delta(\mathbf{r}-\mathbf{r^{\prime}})w_{\mathbf{k}_{\sigma}}(\mathbf{r^{\prime}})w_{\mathbf{l}_{\sigma}}(\mathbf{r})\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{j}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{k}_{\sigma}}\hat{b}_{\mathbf{l}_{\sigma}}$
$=$ $\sum_{\mathbf{i}_{\sigma}}\frac{U_{\mathbf{i}_{\sigma}\mathbf{i}_{\sigma}}}{2}\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{i}_{\sigma}}\hat{b}_{\mathbf{i}_{\sigma}}+\sum_{\left\langle\mathbf{i}_{\sigma},\mathbf{j}_{\sigma}\right\rangle}\frac{U_{\mathbf{i}_{\sigma}\mathbf{j}_{\sigma}}}{2}\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{i}_{\sigma}}\hat{b}_{\mathbf{j}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{j}_{\sigma}}$ (S9)
where
$U_{\mathbf{i}_{\sigma}\mathbf{i}_{\sigma}}\equiv\frac{g_{\sigma\sigma}}{2}\int d^{3}\mathbf{r}\left|w_{\mathbf{i}_{\sigma}}(\mathbf{r})\right|^{4},$ (S10)
and
$U_{\mathbf{i}_{\sigma}\mathbf{j}_{\sigma}}\equiv\frac{g_{\sigma\sigma}}{2}\int d^{3}\mathbf{r}\left|w_{\mathbf{i}_{\sigma}}(\mathbf{r})\right|^{2}\left|w_{\mathbf{j}_{\sigma}}(\mathbf{r})\right|^{2}.$ (S11)
Here $g_{\sigma\sigma}=4\pi\hbar^{2}a_{\sigma\sigma}/\mu$ denotes the intra-species contact interaction strength.
Accordingly, the interaction Hamiltonian between different components takes the form that
$H_{\mathrm{inter}}=\sum_{\langle\mathbf{i}_{\uparrow},\mathbf{j}_{\downarrow}\rangle}U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}\hat{n}_{\mathbf{i}_{\uparrow}}\hat{n}_{\mathbf{j}_{\downarrow}},$ (S12)
where
$U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}\equiv\frac{g_{\uparrow\downarrow}}{2}\int d^{3}\mathbf{r}\left|w_{\mathbf{i}_{\uparrow}}(\mathbf{r})\right|^{2}\left|w_{\mathbf{j}_{\downarrow}}(\mathbf{r})\right|^{2},$ (S13)
with $g_{\uparrow\downarrow}=4\pi\hbar^{2}a_{\uparrow\downarrow}/\mu$ being the inter-species contact interaction strength.
Considering the 2D case, we approximate the Wannier function $w_{\mathbf{i}_{\sigma}}(\mathbf{r})$ as a harmonic oscillator wave function of the lowest energy level, therefore, the Wannier function (at site $\mathbf{0}$ for instance) in the 2D plane reads
$w(x,y)=\sqrt{2\kappa}e^{-\pi\kappa(x^{2}+y^{2})},$ (S14)
where $\kappa\equiv\sqrt{\mu V_{0}/(2\hbar^{2}d^{2})}$ with $V_{0}$ being the lattice depth and $d$ being the lattice constant. As a result, the interlayer interaction strength can be expressed as
$U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}=\frac{g_{\uparrow\downarrow}}{2}\kappa e^{-\pi\kappa\left(\Delta_{x}^{2}+\Delta_{y}^{2}\right)},$ (S15)
where $\Delta_{x}^{2}+\Delta_{y}^{2}\equiv|\mathbf{r_{\mathbf{i}_{\uparrow}}}-\mathbf{r_{\mathbf{j}_{\downarrow}}}|^{2}$, which depends on the twisted angle $\theta$.
Similarly, the strength of the interactions takes the form that
$\begin{cases}U_{\mathbf{i}_{\sigma}\mathbf{i}_{\sigma}}=&\frac{g_{\sigma\sigma}}{2}\kappa\\ U_{\mathbf{i}_{\sigma}\mathbf{j}_{\sigma}}=&\frac{g_{\sigma\sigma}}{2}\kappa e^{-\pi\kappa d^{2}}\\ U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}=&\frac{g_{\uparrow\downarrow}}{2}\kappa e^{-\pi\kappa\left(\Delta_{x}^{2}+\Delta_{y}^{2}\right)}\end{cases}.$ (S16)
In summary, the Hamiltonian of the system reads
$\hat{H}=$ $\sum_{\sigma=\uparrow,\downarrow}\left[-J\sum_{\langle\mathbf{i}_{\sigma},\mathbf{j}_{\sigma}\rangle}\hat{b}_{\mathbf{i_{\sigma}}}^{\dagger}\hat{b}_{\mathbf{j_{\sigma}}}+\sum_{\mathbf{i}_{\sigma}}\frac{U(g_{\sigma\sigma})}{2}\hat{n}_{\mathbf{i}_{\sigma}}(\hat{n}_{\mathbf{i}_{\sigma}}-1)\right]$
$+\sum_{\sigma=\uparrow,\downarrow}\sum_{\langle\mathbf{i}_{\sigma},\mathbf{j}_{\sigma}\rangle}\frac{U_{\mathbf{i}_{\sigma}\mathbf{j}_{\sigma}}(g_{\sigma\sigma})}{2}\hat{n}_{\mathbf{i}_{\sigma}}\hat{n}_{\mathbf{j}_{\sigma}}$
$+\sum_{\langle\mathbf{i}_{\uparrow},\mathbf{j}_{\downarrow}\rangle}U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}(g_{\uparrow\downarrow},\theta)\hat{n}_{\mathbf{i}_{\uparrow}}\hat{n}_{\mathbf{j}_{\downarrow}}.$ (S17)
where $\hat{n}_{\mathbf{i}_{\sigma}}\equiv\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{i}_{\sigma}}$ is the particle number operator that counts the number of atoms of species $\sigma$ on site $\mathbf{i}$ of the $\sigma$-layer and $U\equiv U_{\mathbf{i}_{\sigma}\mathbf{i}_{\sigma}}$.
We investigate the physics of this system by calculating its ground state via the bosonic Gutzwiller variational approach (Krauth et al., 1992; Jaksch et al., 1998; Lanatà et al., 2012). Within the framework of mean-field theory, we use the bosonic Gutzwiller variational approach to investigate the ground state properties of the system, with the variational ground state assumed to be the site-factorized form
$|\mathrm{GW}\rangle=|\phi_{1_{\sigma}}\rangle_{1_{\sigma}}\otimes\ldots\otimes|\phi_{N_{\mathrm{lat},\sigma}}\rangle_{N_{\mathrm{lat},\sigma}}.$ (S18)
Here, $N_{\mathrm{lat},\sigma}$ is the total number of the monolayer lattice sites in $\sigma$-layer and $|\phi_{\mathbf{i}_{\sigma}}\rangle_{\mathbf{\mathbf{i}_{\sigma}}}=\sum_{n=0}^{\infty}c_{n}^{(\mathbf{\mathbf{i}_{\sigma})}}|n\rangle_{\mathbf{\mathbf{i}_{\sigma}}}$ is the local wave function at site $\mathbf{i}$ of $\sigma$-layer with $|n\rangle_{\mathbf{i}_{\sigma}}$ being the corresponding local occupation number state and $c_{n}^{(\mathbf{\mathbf{i}_{\sigma})}}$ being the variational parameter. The ground state is determined by minimizing the total energy of the system within this variational ansatz, i.e.,
$E(\{c_{n}^{(\mathbf{\mathbf{i}_{\sigma})}}\})=$ $\langle\mathrm{GW}|\hat{H}|\mathrm{GW}\rangle$
$=$ $\sum_{\mathbf{i}_{\sigma},\sigma=\uparrow,\downarrow}\left[\frac{U}{2}\frac{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{i}_{\sigma})}\right|^{2}n(n-1)}{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{i}_{\sigma})}\right|^{2}}+\sum_{\mathbf{j}_{\sigma}}\frac{U_{\mathbf{i}_{\sigma}\mathbf{j}_{\sigma}}}{2}\left(\frac{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{i}_{\sigma})}\right|^{2}n}{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{i}_{\sigma})}\right|^{2}}\right)\left(\frac{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{j}_{\sigma})}\right|^{2}n}{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{j}_{\sigma})}\right|^{2}}\right)\right.$
$\left.-J\sum_{\mathbf{j}_{\sigma}}\left(\frac{\sum_{n=0}^{\infty}c_{n}^{*(\mathbf{i}_{\sigma})}c_{n+1}^{(\mathbf{i}_{\sigma})}\sqrt{n+1}}{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{i}_{\sigma})}\right|^{2}}\right)\left(\frac{\sum_{n=0}^{\infty}c_{n}^{(\mathbf{j}_{\sigma})}c_{n+1}^{*(\mathbf{j}_{\sigma})}\sqrt{n+1}}{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{j}_{\sigma})}\right|^{2}}\right)\right]$ (S19)
$+\sum_{\langle\mathbf{i}_{\uparrow},\mathbf{j}_{\downarrow}\rangle}U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}\left(\frac{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{i}_{\uparrow})}\right|^{2}n}{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{i}_{\uparrow})}\right|^{2}}\right)\left(\frac{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{j}_{\downarrow})}\right|^{2}n}{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{j}_{\downarrow})}\right|^{2}}\right).$
During the minimization process, we have tried different initial states to obtain the global energy minimum.
II Phase diagrams for integer filling
Figure S1: Phase diagrams and typical configurations of the MI-MI case and SF-MI case. (a) Phase diagram with respect to the interlayer interaction strength $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ and the hopping amplitude $J/U(g_{\sigma\sigma})$ at the filling with $\rho_{\uparrow}=1$ and $\rho_{\downarrow}=1$ (twist angle is fixed at $\theta=36.87^{\circ}$). (b) Similar to (a), but at the filling with $\rho_{\uparrow}=1/5$ and $\rho_{\downarrow}=1$. (c) and (d) Real-space particle density and superfluid order parameter distributions of the system that correspond to the red dots and the blue squares in (a) and (b), respectively.
In the main text, we have focused on the SF-SF case, where both components have non-integer fillings, rather than on the MI-MI case (both components with integer fillings) or the SF-MI case (one component with a non-integer filling and the other with an integer filling). This choice is due to the large excitation gap caused by the strong interactions in the integer-filling component, which hinders atom tunneling and the formation of rich phases, unlike in the SF-SF case. Here, we present the phase diagrams and typical configurations for the MI-MI and SF-MI cases (see Fig. S1). Similar to the conventional Bose-Hubbard model, when the interlayer interaction strength $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ is small, the components with integer filling favor the homogeneous superfluid phase and Mott-insulator phase [see Figs. S1 (c1) and (c2)], while the component with non-integer filling favor the homogeneous superfluid phase [see Figs. S1 (d1) and (d2)]. In the MI-MI case, even when $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ is large, the density distributions of both components remain almost homogeneous due to the excitation gap [see Figs. S1 (c3) and (c4)]. Conversely, in the SF-MI case, a large $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ drives the component with non-integer filling into an induced Mott-insulator phase [see Fig. S1 (d4)]. Moreover, in the regime of relatively large $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ (where interlayer interactions dominate over intralayer interactions), the intense competition between the integer-filling components due to the interlayer interaction tends to make the system more prone to forming a superfluid phase. However, in regions dominated by interlayer interactions, stronger interlayer interactions make it easier to form a Mott-insulator phase [see Fig. S1 (a)]. In the SF-MI scenario, the high-filling component can create an effective potential that localizes the low-filling component, thereby facilitating the formation of the Mott-insulator phase [see Fig. S1 (b)].

### v4

@@
-Dynamical Moiré Systems in Twisted Bilayer Optical Lattices
+Interaction induced moiré systems in twisted bilayer optical lattices
 Abstract
@@
 Simulation of moiré related physics in cold atom platform is expected to outperform the 2D materials thanks to its advantage of higher tunablility.
-Here, we point out, the cold atom platform enables a new mechanism of moiré lattice formation, with intrinsic “dynamical” character arising from interlayer interaction, in contrast to conventional moiré lattice induced by “static” ways such as single-particle interlayer tunneling or external controls.
+Here, we demonstrate that, the cold atom platform enables a new mechanism of moiré lattice formation, induced by interlayer interaction with intrinsic “dynamical” character, in contrast to conventional moiré lattice induced by “static” ways such as single-particle interlayer tunneling.
 Specifically, we consider a twisted bilayer Bose-Hubbard model with vanishing interlayer tunneling, and the bilayer is solely coupled through interlayer interaction that originates from contact interaction of atoms.
 We find that this system hosts a plethora of novel phases unique to this dynamical lattice, including a variety of Mott insulator (MI) and superfluid (SF) phases either preserving or breaking moiré lattice symmetry, phases with one layer in SF and the other in MI, “interlocked” MI, and self-localized phases at commensurate twist angles, which exhibits the characteristics of Bose glass and quasi-many-body localization in the absence of (quasi)disorder or quasicrystalline lattices.
-Our prediction can be readily observed in current experimental setup of twisted bilayer optical lattices, opening up new avenues for exploring the rich physics of dynamical moiré systems in cold atoms.
-Introduction.—Moiré physics has been extensively studied in the field of condensed matter physics in recent years.
+Our prediction can be readily observed in current experimental setup of twisted bilayer optical lattices, opening up new avenues for exploring the rich physics of interaction induced moiré systems in cold atoms.
+I Introduction
+Moiré physics has been extensively studied in the field of condensed matter physics in recent years.
 One prominent feature that stimulates tremendous interest is the emergence of flat band by simply tuning the twist angle of bilayer two-dimensional (2D) materials, which offers an intriguing platform for exploring phases that electron-electron interaction plays a significant role, such as correlated insulators Cao et al.
@@
 (2021c); Zhou et al.
+(2021); Qiu et al.
 (2023), $\mathbb{Z}_{2}$ topological insulators Wu et al.
@@
 Moreover, being free from lattice relaxation, strain, and disorder, these platforms support pristine twisted bilayer models, allowing direct comparisons between theories and experiments.
-So far, the moiré lattices observed in both 2D materials and cold atoms share a common feature: they are “static”, in the sense that their forms are uniquely fixed by deterministic factors such as interlayer tunneling, structural relaxations, or external controls like dielectric screening Bistritzer and MacDonald (2011); Wang et al.
+So far, the moiré lattices observed in both 2D materials and cold atom studies share a common characteristic: they are “static”, in the sense that their form is uniquely determined by interlayer tunneling, structural relaxations, or external controls such as dielectric screening Bistritzer and MacDonald (2011); Wang et al.
 (2017); Ruiz-Tijerina and Fal’ko (2019); Li et al.
@@
 (2024).
-In this work, we propose a distinct mechanism of realizing moiré lattice solely induced by particle-particle interaction from two layers or pseudospins, which we refer to as “dynamical” moiré lattice  [^1] This is to be contrasted with dynamical control of moiré lattice already realized in 2D materials, where the moiré potential in the target layer is uniquely imprinted by substrate layers and has no feedback on the substrate layers Kim et al.
+Here in this work, we propose a distinct mechanism of realizing moiré lattice purely induced by particle-particle interaction from two layers or two pseudospins, which we refer to as “dynamical” moiré lattice.
+With vanishing interlayer coupling at single-particle level, the two layers are completely isolated, and thus there is no moiré lattice at this level.
+For each layer/pseudospin of lattice, the effect of the other layer is brought by particle-particle interaction, which relies on particle density, a dynamical variable  [^1] Therefore, each layer can feel dynamical feedback from the other layer and the ultimate density distribution is a compromise between the two layers.
+This is to be contrasted with dynamical control of moiré lattice already realized in 2D materials, where the moiré potential in target layer is uniquely imprinted by substrate layers and has no feedback on the substrate layers Kim et al.
 (2024); Zhang et al.
@@
 (2024).
-For each layer/pseudospin of lattice, the effect of the other layer is brought by particle-particle interaction, which relies on particle density distribution, an intrinsic dynamical variable.
-Therefore, each layer can feel dynamical feedback from the other layer and the ultimate density distribution is a compromise between the two layers.
-Realizing such a dynamical moiré system in twisted 2D materials is challenging, but it can be easily achieved in current cold atom platforms by simply turning off interspin coupling in twisted spin-dependent lattices Meng et al.
+Realizing such a dynamical moiré system in twisted 2D materials is challenging, but it can be easily achieved in modern cold atom platforms by simply turning off interspin coupling in the twisted spin-dependent lattices Meng et al.
 (2023).
@@
 The left (right) column of each subplot corresponds to the ($\uparrow$($\downarrow$)↑ ( ↓ )-layer, respectively.
-From left to right, up to down, ($\theta$, $\rho_{\uparrow}$, $\rho_{\downarrow}$, $g_{\uparrow\downarrow}/g_{\sigma\sigma}$, $J/U(g_{\sigma\sigma})$) assume the values ($7.63^{\circ}$, $3/5$, $3/5$, $0.5$, $0.06$), ($36.87^{\circ}$, $4/5$, $4/5$, $0.9$, $0.048$), ($36.87^{\circ}$, $7/10$, $4/5$, $0.9$, $0.04$), ($22.62^{\circ}$, $2/5$, $3/5$, $0.9$, $0.024$), ($36.87^{\circ}$, $4/5$, $4/5$, $0.3$, $0.008$), and ($7.63^{\circ}$, $7/10$, $3/5$, $0.5$, $0.012$), respectively.
-See text for more details.
-Specifically, we consider two species of ultracold atoms loaded in a twisted bilayer spin-dependent optical lattice and find that, an emergent moiré lattice from particle-interaction can form, hosting much more rich phases than static lattice counterpart (see Fig.
+From left to right, up to down, ($\theta$, $\rho_{\uparrow}$, $\rho_{\downarrow}$, $g_{\uparrow\downarrow}/g_{\sigma\sigma}$, $J/U(g_{\sigma\sigma})$) assume the values ($7.63^{\circ}$, $3/5$, $3/5$, $0.5$, $0.06$), ($36.87^{\circ}$, $4/5$, $4/5$, $0.9$, $0.048$), ($36.87^{\circ}$, $7/10$, $4/5$, $0.9$, $0.04$), ($22.62^{\circ}$, $2/5$, $3/5$, $0.9$, $0.024$), ($36.87^{\circ}$, $4/5$, $4/5$, $0.3$, $0.008$), and ($7.63^{\circ}$, $2/5$, $3/5$, $0.5$, $0.024$), respectively.
+Specifically, we consider two species of ultracold atoms loaded in a twisted bilayer optical lattice and find that, an emergent moiré lattice from particle interaction can form, hosting much more rich phases than static lattice counterpart (see Fig.
 1), including phases with and without moiré lattice symmetry and Mott insulator (MI) phases induced by interlayer interaction even for partial filling.
 More interestingly, for total unit filling, the ground state can become an “interlocked” MI with spontaneously formed moiré lattice.
-For a general total filling, the Bose glass phase exists, suggesting that the ground state tends to localize due to the interplay between twisting and interlayer interaction, even at commensurate twist angles.
-System and model.—Motivated by the recent realization of the twisted bilayer optical lattices Meng et al.
-(2023), we consider a closely related but simpler experimental set-up, namely, a twisted bilayer optical lattice system without the additional microwave field for realizing the interlayer coupling in the recent experiments Meng et al.
+For a general total filling, the Bose glass (BG) phase exists, suggesting that the ground state tends to localize due to the interplay between twisting and interlayer interaction, even at commensurate twist angles.
+II System and model
+Motivated by the recent realization of the twisted bilayer optical lattices Meng et al.
+(2023), we consider a closely related but simpler experimental setup, namely, a twisted bilayer optical lattice system without the additional microwave field for realizing the interlayer coupling in the recent experiments Meng et al.
 (2023).
@@
 1(a)].
-The physics of this system at low filling can be described by a Bose-Hubbard type model within the lowest band approximation (see Supplemental Material (SM) Sup for a detailed derivation),
+The physics of this system at low filling can be described by a Bose-Hubbard type model within the lowest band approximation [see Appendix A for a detailed derivation],
 $\hat{H}=$ $\sum_{\sigma=\uparrow,\downarrow}\left[-J\sum_{\langle\mathbf{i}_{\sigma},\mathbf{j}_{\sigma}\rangle}\hat{b}_{\mathbf{i_{\sigma}}}^{\dagger}\hat{b}_{\mathbf{j_{\sigma}}}+\sum_{\mathbf{i}_{\sigma}}\frac{U(g_{\sigma\sigma})}{2}\hat{n}_{\mathbf{i}_{\sigma}}(\hat{n}_{\mathbf{i}_{\sigma}}-1)\right]$
@@
 1(c)], and consists of interlayer interaction terms $U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}(g_{\uparrow\downarrow},\theta)\hat{n}_{\mathbf{i}_{\uparrow}}\hat{n}_{\mathbf{j}_{\downarrow}}$, with the interaction strength $U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}(g_{\uparrow\downarrow},\theta)$ assuming the twist angle $\theta$ dependence.
-This dependence reflects how the overlaps between Wannier wave functions of different sites on different layers are changed when one tunes $\theta$.
-Figure 2: Phase diagrams and typical configurations at the balanced and imbalanced fillings.
-(a) Phase diagram with respect to the interlayer interaction strength $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ and the hopping amplitude $J/U(g_{\sigma\sigma})$ at the balanced filling with $\rho_{\uparrow}=4/5$ and $\rho_{\downarrow}=4/5$ (twist angle is fixed at $\theta=36.87^{\circ}$).
-(b) Similar to (a), but at the imbalanced filling with $\rho_{\uparrow}=1/5$ and $\rho_{\downarrow}=4/5$.
-(c) and (d) Real-space particle density and superfluid order parameter distributions of the system that correspond to the red dots and the blue squares in (a) and (b), respectively.
-See text for more details.
+As the overlaps of the wave functions depend directly on the spatial distance of sites $\mathbf{i}_{\uparrow}$ and $\mathbf{j}_{\downarrow}$, the interaction strength $U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}(g_{\uparrow\downarrow},\theta)$ has explicit dependence on both spatial distance and twist angle $\theta$.
+While many studies have investigated multicomponent Bose-Hubbard models in the context of ultracold atoms, our study focuses on the largely unexplored many-body physics of two-component ultracold bosons in twisted bilayer optical lattices.
+The rich physics arising from the twist-dependent interlayer interaction $U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}(g_{\uparrow\downarrow},\theta)$ in such systems has not yet been systematically studied.
 We investigate this system by calculating its ground state via the bosonic Gutzwiller variational approach Krauth et al.
@@
 (1998); Lanatà et al.
-(2012) (see SM Sup for technical details).
+(2012) (see Appendix A for technical details).
 We have chosen the twist angle $\theta=2\arctan(\bar{m}/\bar{n})$ with $\bar{m},\bar{n}$ being two natural numbers Wang et al.
@@
 (2023), i.e., the commensurate condition for twisted square lattice, and focus on the case with $g_{\uparrow\uparrow}=g_{\downarrow\downarrow}$.
-If not specified otherwise, we perform the numerical simulation on a $37\times 37$ twisted bilayer optical lattice with an external 2D cylinder box trap Navon et al.
+We perform the numerical simulation on a $37\times 37$ twisted bilayer optical lattice with an external 2D cylinder box trap Navon et al.
 (2021) whose radius is 18 lattice constants (denoted by the solid circles in the figures).
-In the following, we focus on the most interesting case of fractional filling for both layers, and reserve the results of integer filling in the Supplemental Material Sup (the impact of interlayer interaction is not as dramatic in the latter case).
+We additionally investigate the finite-size effects and find no significant differences between the results here and those on a larger system (see the Appendix C).
+In the following, we focus on the most interesting case of fractional filling for both layers, and reserve the results of integer filling in the Appendix B (the impact of interlayer interaction is not as dramatic in the latter case).
 Before presenting the numerical results, we first notice that in conventional moiré systems Bistritzer and MacDonald (2011); Wang et al.
 (2017); Ruiz-Tijerina and Fal’ko (2019); Meng et al.
-(2023), one layer of lattice actually influences the physics on the other layer in a “static” way, in the sense that the influences are solely determined by the configuration of the twisted bilayer lattice through the single-particle interlayer tunneling.
-In sharp contrast, here, these influences are “dynamical”, in the sense that one layer of lattice impose its influences on the physics of the other via the dynamical variables on this layer (particle number operators in this case), as clearly manifested by the interlayer interaction terms of the Hamiltonian (1), giving rise to a new type of moiré system we dubbed as “dynamical moiré system”.
-This intrinsic dynamical character could make the system feature not only similar phases in the conventional moiré system, but also many intriguing phases without counterpart in the latter [see Fig.
+(2023); Qiu et al.
+(2023), the moiré lattices in twisted bilayers are induced by single-particle interlayer tunneling.
+Particle interactions can be further considered within the framework of these preformed moiré lattices.
+In sharp contrast, here, the twisted bilayers are completely isolated at single-particle level, and the moiré lattice arises purely from the interlayer interaction terms of the Hamiltonian (1).
+This novel mechanism of moiré formation could make the system feature not only similar phases in the conventional moiré system, but also many intriguing phases without counterpart in the latter [see Fig.
 1(d) for an overview].
-Conventional moiré physics without interlayer tunneling.—In conventional bosonic moiré systems, the interlayer tunneling causes layer hybridization of single particle orbitals, and thus particle filling on the hybridized orbitals results in phases with two layer sharing the same character.
+Figure 2: Phase diagrams and typical configurations at the balanced and imbalanced fillings.
+(a) Phase diagram with respect to the interlayer interaction strength $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ and the hopping amplitude $J/U(g_{\sigma\sigma})$ at the balanced filling with $\rho_{\uparrow}=4/5$ and $\rho_{\downarrow}=4/5$ (twist angle is fixed at $\theta=36.87^{\circ}$).
+The magenta region of the phase diagram contains the BG phase.
+(b) Similar to (a), but at the imbalanced filling with $\rho_{\uparrow}=1/5$ and $\rho_{\downarrow}=4/5$.
+(c) and (d) Real-space particle density and superfluid order parameter distributions of the system that correspond to the red dots and the blue squares in (a) and (b), respectively.
+III Results
+III.1 Conventional moiré physics without interlayer tunneling
+In conventional bosonic moiré systems, the interlayer tunneling causes layer hybridization of single particle orbitals, and thus particle filling on the hybridized orbitals results in phases with two layer sharing the same character.
 Indeed, as observed in recent experiments Meng et al.
@@
 2(a), we map out the phase diagram of the system in a balanced filling case with $\rho_{\uparrow}=\rho_{\downarrow}=4/5$ ($\rho_{\sigma}\equiv\langle\sum_{\mathbf{i}_{\sigma}}\hat{n}_{\mathbf{i}_{\sigma}}\rangle/N_{\sigma}^{\mathrm{lat}}$ with $N_{\sigma}^{\mathrm{lat}}$ being the number of lattice sites of the $\sigma$-layer within the cylinder trap).
-In the relatively large hopping regime, we can see that the system indeed support SF-SF phase (denoted as SF in the phase diagrams) similar to conventional static moiré systems, with both layers manifesting the SF order parameter $\phi_{\mathbf{i}_{\sigma}}\equiv\langle\hat{b}_{\mathbf{i}_{\sigma}}\rangle$ and density distribution $\langle\hat{n}_{\mathbf{i}_{\sigma}}\rangle$ assuming the moiré lattice symmetry [see the first plot in Fig.
-2(c)].
-While in the relatively small hopping regime, we notice that despite the filling factor on each layer being non-integer, increasing the inter-species interaction strength $g_{\uparrow\downarrow}$ can drive the system into the induced MI phase characterized by regions with integer-filled particles and vanishing SF order parameter in either layer [see the last plot in Fig.
-2(c)], including MI-SF, MI-MI and SF-MI phases.
+In the relatively large hopping regime, we can see that the system indeed support SF-SF phase (denoted as SF in the phase diagrams) similar to conventional static moiré systems, with both layers manifesting the SF order parameter $\phi_{\mathbf{i}_{\sigma}}\equiv\langle\hat{b}_{\mathbf{i}_{\sigma}}\rangle$ and density distribution $\langle\hat{n}_{\mathbf{i}_{\sigma}}\rangle$ assuming the moiré lattice symmetry [see Fig.
+2(c1)].
+While in the relatively small hopping regime, we notice that despite the filling factor on each layer being non-integer, increasing the inter-species interaction strength $g_{\uparrow\downarrow}$ can drive the system into the induced MI phase characterized by regions with integer-filled particles and vanishing SF order parameter in either layer [see Fig.
+2(c4)], including MI-SF, MI-MI and SF-MI phases.
 This is in sharp contrast to the single-component Bose gases in optical lattices, where the MI can only exist at integer fillings.
@@
 However, it’s important to recognize a key difference here: the effective potential is a dynamical one determined by the dynamical variables, and hence does not necessarily share the same spatial symmetry of the underlying optical lattice.
-For balanced filling case, indeed, we have not found MI phase with both layer assumes the moiré lattice symmetry.
-Instead, as shown in the second and third plot in Fig.
-2(c), we find this dynamical moiré system can manifest rich discrete rotational symmetric MI and SF phases with no counterparts in the conventional bosonic moiré systems.
+For balanced filling case, indeed, we have not found MI phase with both layer assuming the moiré lattice symmetry.
+Instead, as shown in the Figs.
+2(c2) and (c3), we find this dynamical moiré system can manifest rich discrete rotational symmetric MI and SF phases with no counterparts in the conventional bosonic moiré systems.
 However, away from the balanced filling, we indeed find the system can support MI phase with density distributions on both layers assuming the moiré lattice symmetry.
-For instance, as shown in the last two plots in Fig.
-2(d), this kind of moiré MI phase can emerge in the induced MI regime of the system at an imbalanced filling $\rho_{\uparrow}=1/5,\rho_{\downarrow}=4/5$.
+For instance, as shown in the Fig.
+2(d4), this kind of moiré MI phase can emerge in the induced MI regime of the system at an imbalanced filling $\rho_{\uparrow}=1/5,\rho_{\downarrow}=4/5$.
+In this case, $\uparrow$-particles occupy overlapping sites while $\downarrow$-particles occupy non-overlapping sites within each moiré period [see the Fig.
+1(c)], minimizing interlayer repulsion and preserving lattice symmetry.
+If the filling factors deviate from $(\rho_{\uparrow}=1/5,\rho_{\downarrow}=4/5)$ for $\theta=36.87^{\circ}$, the remaining particles or vacant sites are likely to break the moiré lattice symmetry.
+Similarly, the system with other twist angle [e.g.
+the $(\theta=22.62^{\circ},\rho_{\uparrow}=2/5,\rho_{\downarrow}=3/5)$ case shown in the lower left plot of Fig.
+1(d)] also hosts this kind of “interlocked” MI with moiré lattice symmetry.
+By contrast, for a SF phase with moiré lattice symmetry, the filling factor requirements are less restrictive because the number of atoms per site is not constrained to integer values, allowing particles to distribute themselves in accordance with the moiré lattice symmetry.
+We also notice that the interplay between interlayer repulsion and filling imbalance gives rise to the difference in the critical $J/U(g_{\sigma\sigma})$ shown in Figs.
+2(a) and (b): in the imbalanced case ($\rho_{\uparrow}=1/5$), the larger interparticle distance on the $\uparrow$-layer requires a higher $J/U(g_{\sigma\sigma})$ for coherence and transition into the SF phase compared to the balanced case ($\rho_{\downarrow}=4/5$).
 Although the dynamical moiré system can manifest SF or MI phase similar to the ones found in conventional bosonic moiré systems Meng et al.
 (2023), there are fundamental differences.
-Instead of being coupled by interlayer tunneling, the atoms on different layers are coupled by the interlayer density-density interactions.
-Consequentially, the total particle number of each layer is conserved and can be different.
+Instead of being coupled by interlayer tunneling, the atoms on different layers are coupled by the interlayer density-density interactions, which preserve the independent U(1) symmetries in each layer.
+Consequently, the total particle number of each layer is conserved and can be different.
 These fundamental differences thus naturally indicate much richer physics with no counterparts in conventional bosonic moiré systems can emerge, as we shall discuss below.
-Beyond the conventional moiré physics.—By comparing the physics at balanced and imbalanced filling as showcased in Fig.
-2, one clearly notices the strong physical influences of the filling difference.
-Since the total particle number of each layer is conserved separately, the filling difference between the two layers in fact provides an extra tuning knob of new physics compared with the conventional bosonic moiré systems.
+III.2 Beyond the conventional moiré physics
+Since the total particle number of each layer is conserved separately, the filling difference between the two layers in fact provides an extra tuning knob of new physics compared with the conventional bosonic moiré systems, as showcased in Fig.
+2.
 In Fig.
-3(a), we systematically investigate the influences of the filling difference by keeping the filling factor of the $\downarrow$-layer fixed at $\rho_{\downarrow}=4/5$ and changing the filling factor of the $\uparrow$-layer $\rho_{\uparrow}$ from $1/10$ all the way to $4/5$.
-As we can see from the first plot in Fig.
-3(c), when the filling of the $\uparrow$-layer is small ($\rho_{\uparrow}=1/10$), the $\downarrow$-layer is in a SF phase and it imposes strong influences on atoms on the $\uparrow$-layer by localizing them and making them in induced MI phase despite $\rho_{\uparrow}=1/10$ being non-integer.
-As the $\rho_{\uparrow}$ gradually increases [see the last two plots in Fig.
-3(c)], the $\uparrow$-layer also imposes an enhanced localization effects on the atoms on the $\downarrow$-layer by driving a SF-MI transition on this layer.
+3(a), we systematically investigate the influences of the filling difference by keeping the filling factor of the $\downarrow$-layer fixed at $\rho_{\downarrow}=4/5$ and changing the filling factor of the $\uparrow$-layer $\rho_{\uparrow}$ from $1/10$ to $4/5$.
+As we can see from the Fig.
+3(c1), when the filling of the $\uparrow$-layer is small ($\rho_{\uparrow}=1/10$), the $\downarrow$-layer is in a SF phase and it imposes strong influences on atoms on the $\uparrow$-layer by localizing them and making them in induced MI phase despite $\rho_{\uparrow}=1/10$ being non-integer.
+As the $\rho_{\uparrow}$ gradually increases [see Figs.
+3(c2) and (c3)], the $\uparrow$-layer also imposes an enhanced localization effects on the atoms on the $\downarrow$-layer by driving a SF-MI transition on this layer.
 We also observe that the average localization effect experienced by each atom in the $\uparrow$-layer decreases, making the $\uparrow$-layer transition into a SF phase at relatively high filling with $\rho_{\uparrow}=1/2$.
@@
 (2015).
-More interestingly, for the total unit filling, the system hosts a kind of “interlocked” MI with “interlocked” density distributions [see the second plot of Fig.
-3(c), the lower left plot of Fig.
-1(d) and the fourth plot of Fig.
-2(d)].
 Similar to the conventional moiré system, the twist angle also has a significant impact here, as clearly shown in Fig.
@@
 For the case at the filling $(\rho_{\uparrow}=1/10,\rho_{\downarrow}=3/5)$, in the intermediate hopping regime, we see that at a relatively small twist angle $\theta=7.63^{\circ}$, the system is in the SF-SF phase.
-As the twist angle increases to $\theta=22.62^{\circ}$, the atoms in the $\uparrow$-layer are localized due to the interaction from the $\downarrow$-layer and enter the induced MI phase, forming the phase MI-SF [see the second plot of Fig.
-3(d)].
-Similar situation is also found at even larger twist angle, for instance, $\theta=36.87^{\circ}$ [see the last plot of Fig.
-3(d)].
-Interestingly, we notice that density distributions in the last two plots of Fig.
-3(d) assume rather random pattern.
+As the twist angle increases to $\theta=22.62^{\circ}$, the atoms in the $\uparrow$-layer are localized due to the interaction from the $\downarrow$-layer and enter the induced MI phase, forming the phase MI-SF [see Fig.
+3(d2)].
+Similar situation is also found at even larger twist angle, for instance, $\theta=36.87^{\circ}$ [see Fig.
+3(d3)].
+Interestingly, we notice that density distributions in the Figs.
+3(d2) and (d3) assume rather random pattern.
 In these cases, the atoms on $\downarrow$-layer in fact provide a “disordered” effective potential to the atoms lying on the $\uparrow$-layer.
 Thus, one naturally expects that this “disordered” effective potential quenches hopping of the $\uparrow$-atoms and makes them localized, which can be interpreted as the interlayer interaction induced “self-localization” between the atoms on the two layers.
-Among the states with irregular density pattern, such as the third plot of Fig.
-3(d) and the last plot of Fig.
-1(d), there exist states with disconnected superfluid islands, exhibiting the characteristics of Bose glass.
-Thus, this system supports Bose glass without extrinsic disorder, in contrast to previous studies with either (quasi)disorder or quasicrystalline lattices Fisher et al.
+The presence of the “disordered” effective potential is associated with phases with the spontaneous breaking of $C_{4}$ rotational symmetry of the system [see Fig.
+3(c1) and Fig.
+3(d)], indicating that there are four degenerate ground states (see the Appendix D for the degenerate ground states).
+Figure 3: (a) Phase diagram with respect to $\rho_{\uparrow}$ and $J/U(g_{\sigma\sigma})$ with other parameters fixed at $g_{\uparrow\downarrow}/g_{\sigma\sigma}=0.9$, $\theta=36.87^{\circ}$, and $\rho_{\downarrow}=4/5$.
+(b) Phase diagram with respect to the twist angle $\theta$ and $J/U(g_{\sigma\sigma})$ with other parameters fixed at $g_{\uparrow\downarrow}/g_{\sigma\sigma}=0.9$, $\rho_{\uparrow}=1/10$, and $\rho_{\downarrow}=3/5$.
+The magenta regions of the phase diagrams contain the BG phase.
+(c) and (d) Real-space particle density and superfluid order parameter distributions of the system that correspond to the red dots and the blue squares in (a) and (b), respectively.
+Among the states with irregular density pattern, such as the Fig.
+3(d3) and the last plot of Fig.
+1(d), there exist states with disconnected superfluid islands, exhibiting the characteristics of BG (see the Appendix E for exact identification).
+Thus, this system supports BG without extrinsic disorder (see the magenta region corresponding to the BG phase in the phase diagrams shown in Figs.
+2 and 3), in contrast to previous studies with either (quasi)disorder or quasicrystalline lattices Fisher et al.
 (1989); Fallani et al.
@@
 (2016); Johnstone et al.
-(2021); Ciardi et al.
+(2021a); Ciardi et al.
 (2023).
-The presence of Bose glass is associated with the spontaneous breaking of $C_{4}$ rotational symmetry of the system, indicating that there are four degenerate ground states.
-As the presence of Bose glass at ground state usually implies many-body localization (MBL) at excited states Abanin et al.
+BG phase in our system is characterized by the presence of isolated, non-percolated superfluid islands, and usually surrounded by MI regions.
+Therefore, in the phase diagrams, the BG phase is located within the broader “induced MI” region, except the filling-angle-match case [see the Fig.
+2 (b)] whose density distributions with the moiré lattice symmetry would not provide an effective disordered potential.
+Consistent with previous studies, we have found many metastable configurations associated with the BG phase in our system (see the Appendix D for the metastable configurations).
+As the presence of BG at ground state usually implies many-body localization (MBL) at excited states Abanin et al.
 (2019), this system also probably hosts (quasi-)MBL without extrinsic disorder Schiulaz et al.
@@
 (2021).
-Note that here either the Bose glass or the possible quasi-MBL are found at commensurate twist angles, i.e., the system has translational invariance in the bulk, in contrast to previously investigated localization at incommensurate angles Huang and Liu (2019); Paul et al.
+Note that here either the BG or the possible quasi-MBL are found at commensurate twist angles, i.e., the system has translational invariance in the bulk, in contrast to previously investigated localization at incommensurate angles Huang and Liu (2019); Paul et al.
 (2023); Madroñero et al.
@@
 Further exploration of the relation between the localization found here with MBL and the generalization to incommensurate twist angles will be an intriguing subject for future study.
-Figure 3: (a) Phase diagram with respect to $\rho_{\uparrow}$ and $J/U(g_{\sigma\sigma})$ with other parameters fixed at $g_{\uparrow\downarrow}/g_{\sigma\sigma}=0.9$, $\theta=36.87^{\circ}$, and $\rho_{\downarrow}=4/5$.
-(b) Phase diagram with respect to the twist angle $\theta$ and $J/U(g_{\sigma\sigma})$ with other parameters fixed at $g_{\uparrow\downarrow}/g_{\sigma\sigma}=0.9$, $\rho_{\uparrow}=1/10$, and $\rho_{\downarrow}=3/5$.
-(c) and (d) Real-space particle density and superfluid order parameter distributions of the system that correspond to the red dots and the blue squares in (a) and (b), respectively.
-See text for more details.
-In light of the tremendous experimental progress in ultracold atomic physics, the rich physics in this dynamical moiré system can be readily observed using current experimental setups Meng et al.
+In light of the tremendous experimental advancements in ultracold atomic physics, we believe that the rich physics in this dynamical moiré system can be readily observed using current experimental setups Meng et al.
 (2023).
-Specifically, $J/U(g_{\sigma\sigma})$ can be tuned by adjusting the depth of optical lattice potential.
+More specifically, the $J/U(g_{\sigma\sigma})$ can be tuned by adjusting the depth of optical lattice potential.
 The interlayer interaction $U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}(g_{\uparrow\downarrow},\theta)$ can be altered by either tuning the vertical displacement between the two spin-dependent lattices, or by utilizing the Feshbach resonance  [^2] The Feshbach resonance technique enables the direct tuning of $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ in the phase diagrams.
@@
 The exotic SF, MI phases, and various self-localization behavior can be directly observed via the quantum gas microscope Gross and Bakr (2021).
-Conclusions.—We propose a new type of moiré system in twisted bilayer optical lattices, namely, the dynamical moiré system that hosts a wide range of unique physical phenomena.
-It can not only accommodate moiré SF phase similar to the one found in conventional static bosonic moiré system, but also give rise to rich physics without static counterparts.
-These include “interlocked” MI, exotic SF and MI phases breaking moiré lattice symmetry, phases with one layer in SF and the other in MI, localization behavior similar to Bose glass and quasi-MBL without extrinsic disorder.
-Since both the dynamical moiré system itself and its experimental realization is even simpler than the conventional bosonic moiré system realized in experiments recently Meng et al.
-(2023), we believe our work will stimulate both further theoretical research on this type of dynamical moiré system, especially the possible intrinsic connection between the self-localization found here and quasi-MBL, generalization to fermionic systems, as well as experimental effort in searching for these exotic quantum phases.
+IV Conclusions
+We have proposed a dynamical moiré system induced purely by interlayer particle interaction in twisted bilayer optical lattices, which hosts a wide range of unique physical phenomena.
+both the dynamical moiré system itself and its experimental realization is even simpler than the conventional bosonic moiré system realized in experiments recently Meng et al.
+(2023), we believe our work will stimulate further theoretical research on this type of dynamical moiré system, especially the possible intrinsic connection between the self-localization found here and quasi-MBL, generalization to fermionic systems, as well as experimental effort in searching for these exotic quantum phases.
+In addition, while not present in our purely repulsive interlayer system, other exotic quantum phases like pair superfluid and dimerized phases could exist in bilayer systems with attractive interlayer interactions Macia et al.
+(2014); Cinti et al.
+(2017), warranting further study.
 Acknowledgements.
@@
 2020B1212060066), and START grant of South China Normal University.
-Supplemental Material for “Dynamical Moiré Systems in Twisted Bilayer Optical Lattices”
-I System and model
-Motivated by the recent realization of the twisted bilayer optical lattices (Meng et al., 2023), we consider a closely related but simpler experimental set-up, namely, a twisted bilayer optical lattice system without the additional microwave field that has been used to realize the interlayer coupling in the recent experiments (Meng et al., 2023).
-More specifically, the system under study consists of two species of bosonic atoms that correspond to two hyperfine states (for instance, ${}^{87}\mathrm{Rb}$ atoms in hyperfine states $|F=1,m_{F}=1\rangle$ and $|F=2,m_{F}=0\rangle$ considered in Ref.
-(Meng et al., 2023)), with each species loaded in one layer of a state-dependent twisted bilayer square optical lattice, respectively [see Fig.
-1(a) of the main text for a schematic illustration].
-The physics of this system at low filling can be described by a Bose-Hubbard type model within the lowest band approximation.
-To derive the model Hamiltonian, we start with the following many-body Hamiltonian
-$\hat{H}=$ $\sum_{\sigma=\uparrow,\downarrow}\int d^{3}\mathbf{r}\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r})\left(\frac{-\hbar^{2}}{2\mu}\nabla^{2}+V(\mathrm{\bm{r}})\right)\hat{\psi_{\sigma}}(\mathbf{r})$
-$+\sum_{\sigma=\uparrow,\downarrow}\int d^{3}\mathbf{r}\int d^{3}\mathbf{r}^{\prime}\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r})\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r^{\prime}})\frac{V_{\sigma\sigma}(\mathbf{r}-\mathbf{r^{\prime}})}{2}\hat{\psi}_{\sigma}(\mathbf{r^{\prime}})\hat{\psi}_{\sigma}(\mathbf{r})$ (S1)
+Appendix A Model Hamiltonian and Gutzwiller method
+The low physics of this system can be described by a Bose-Hubbard type model within the lowest band approximation. To derive the model Hamiltonian, we start with the following many-body Hamiltonian
+$\hat{H}=$ $\sum_{\sigma=\uparrow,\downarrow}\int d^{3}\mathbf{r}\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r})\left(\frac{-\hbar^{2}}{2\mu}\nabla^{2}+V(\mathrm{\bm{r}})\right)\hat{\psi_{\sigma}}(\mathbf{r})$ (2)
+$+\sum_{\sigma=\uparrow,\downarrow}\int d^{3}\mathbf{r}\int d^{3}\mathbf{r}^{\prime}\left[\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r})\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r^{\prime}})\frac{V_{\sigma\sigma}(\mathbf{r}-\mathbf{r^{\prime}})}{2}\right.$
+$\left.\hat{\psi}_{\sigma}(\mathbf{r^{\prime}})\hat{\psi}_{\sigma}(\mathbf{r})\right]$
 $+\int d^{3}\mathbf{r}\int d^{3}\bm{\mathbf{r}}^{\prime}\hat{\psi}_{\uparrow}^{\dagger}(\mathbf{r})\hat{\psi}_{\downarrow}^{\dagger}(\mathbf{r^{\prime}})V_{\uparrow\downarrow}(\mathbf{r}-\mathbf{r^{\prime}})\hat{\psi}_{\downarrow}(\mathbf{r^{\prime}})\hat{\psi}_{\uparrow}(\mathbf{r}),$
@@
 $V_{\sigma\sigma}(\mathbf{r}-\mathbf{r^{\prime}})$ $=\frac{4\pi\hbar^{2}a_{\sigma\sigma}}{\mu}\delta(\mathbf{r}-\mathbf{r^{\prime}}),$
-$V_{\uparrow\downarrow}(\mathbf{r}-\mathbf{r^{\prime}})$ $=\frac{4\pi\hbar^{2}a_{\uparrow\downarrow}}{\mu}\delta(\mathbf{r}-\mathbf{r^{\prime}}),$ (S2)
+$V_{\uparrow\downarrow}(\mathbf{r}-\mathbf{r^{\prime}})$ $=\frac{4\pi\hbar^{2}a_{\uparrow\downarrow}}{\mu}\delta(\mathbf{r}-\mathbf{r^{\prime}}),$ (3)
 with $a_{\sigma\sigma}$ ($a_{\uparrow\downarrow}$) being the intra(inter)-species $s$-wave scattering length.
-The first term of the Hamiltonian (S1) is the single particle term, and the second term describes the intra-species interaction of atoms in the same lattice layer, and the third term describes the inter-species interaction between atoms of different lattice layer.
+The first term of the Hamiltonian (2) is the single particle term, and the second term describes the intra-species interaction of atoms in the same lattice layer, and the third term describes the inter-species interaction between atoms of different lattice layer.
 Within the second quantization formalism, we can define the creation operator for the lowest band Wannier orbital $w_{\mathbf{i}_{\sigma}}(\mathbf{r})=\langle\mathbf{r}\mid\mathbf{i}_{\sigma}\rangle$ at site $\mathbf{i}$ of $\sigma$-layer,
-$\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}=\int d^{3}\mathbf{r}w_{\mathbf{i}_{\sigma}}(\mathbf{r})\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r}).$ (S3)
-Therefore,
-$\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r})=\sum_{\mathbf{i}_{\sigma}}\langle\mathbf{i}_{\sigma}\mid\mathbf{r}\rangle\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}=\sum_{\mathbf{i}_{\sigma}}w_{\mathbf{i}_{\sigma}}^{*}(\mathbf{r})\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger},$ (S4)
-$\hat{\psi}_{\sigma}(\mathbf{r})=\sum_{\mathbf{i}_{\sigma}}\langle\mathbf{r}\mid\mathbf{i}_{\sigma}\rangle\hat{b}_{\mathbf{i}_{\sigma}}=\sum_{\mathbf{i}_{\sigma}}w_{\mathbf{i}_{\sigma}}(\mathbf{r})\hat{b}_{\mathbf{i}_{\sigma}},$ (S5)
-where $\hat{b}_{\mathbf{i_{\sigma}}}^{\dagger}(\hat{b}_{\mathbf{i}_{\sigma}})$ is the creation (annihilation) operator at site $\mathbf{i}$ of the lattice layer $\sigma$ in the Wannier basis.
+$\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}=\int d^{3}\mathbf{r}w_{\mathbf{i}_{\sigma}}(\mathbf{r})\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r}).$ (4)
 For a relatively shallow optical lattice, the single particle Hamiltonian within the tight-binding approximation restricted to lowest band reads
-$H_{\mathrm{s.p.},\sigma}=$ $\int d^{3}\mathbf{r}\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r})\left(\frac{-\hbar^{2}}{2\mu}\nabla^{2}+V(\mathrm{\bm{r}})\right)\hat{\psi_{\sigma}}(\mathbf{r})$
-$=$ $\sum_{\mathbf{i}_{\sigma},\mathbf{j}_{\sigma}}\int d^{3}\mathbf{r}w_{\mathbf{i}_{\sigma}}^{*}(\mathbf{r})\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\left(\frac{-\hbar^{2}}{2\mu}\nabla^{2}+V(\mathrm{\bm{r}})\right)w_{\mathbf{j}_{\sigma}}(\mathbf{r})\hat{b}_{\mathbf{j}_{\sigma}}$ (S6)
+$H_{\mathrm{s.p.},\sigma}=$ $\int d^{3}\mathbf{r}\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r})\left(\frac{-\hbar^{2}}{2\mu}\nabla^{2}+V(\mathrm{\bm{r}})\right)\hat{\psi_{\sigma}}(\mathbf{r})$ (5)
+$=$ $\sum_{\mathbf{i}_{\sigma},\mathbf{j}_{\sigma}}\int d^{3}\mathbf{r}w_{\mathbf{i}_{\sigma}}^{*}(\mathbf{r})\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\left(\frac{-\hbar^{2}}{2\mu}\nabla^{2}+V(\mathrm{\bm{r}})\right)w_{\mathbf{j}_{\sigma}}(\mathbf{r})\hat{b}_{\mathbf{j}_{\sigma}}$
 $=$ $\sum_{\mathbf{i}_{\sigma}}\epsilon_{\mathbf{i}_{\sigma}}\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{i}_{\sigma}}-J\sum_{\left\langle\mathbf{i}_{\sigma},\mathbf{j}_{\sigma}\right\rangle}\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{j}_{\sigma}},$
 […]
-$\epsilon_{\mathbf{i}_{\sigma}}\equiv\int d^{3}\mathbf{r}w_{\mathbf{i}_{\sigma}}^{*}(\mathbf{r})[\frac{-\hbar^{2}}{2\mu}\nabla^{2}+V(\mathrm{\bm{r}})]w_{\mathbf{i}_{\sigma}}(\mathbf{r}),$ (S7)
+$\epsilon_{\mathbf{i}_{\sigma}}\equiv\int d^{3}\mathbf{r}w_{\mathbf{i}_{\sigma}}^{*}(\mathbf{r})[\frac{-\hbar^{2}}{2\mu}\nabla^{2}+V(\mathrm{\bm{r}})]w_{\mathbf{i}_{\sigma}}(\mathbf{r}),$ (6)
 and
-$J\equiv-\int d^{3}\mathbf{r}w_{\mathbf{i}_{\sigma}}^{*}(\mathbf{r})[\frac{-\hbar^{2}}{2\mu}\nabla^{2}+V(\mathrm{\bm{r}})]w_{\mathbf{j}_{\sigma}}(\mathbf{r}).$ (S8)
+$J\equiv-\int d^{3}\mathbf{r}w_{\mathbf{i}_{\sigma}}^{*}(\mathbf{r})[\frac{-\hbar^{2}}{2\mu}\nabla^{2}+V(\mathrm{\bm{r}})]w_{\mathbf{j}_{\sigma}}(\mathbf{r}).$ (7)
 For a sufficiently dilute, weakly interacting atomic gas, the short ranged interactions covering the nearest neighbor sites of single component are taken into consideration.
 The interaction Hamiltonian of single component reads
-$H_{\mathrm{intra},\sigma}=$ $\int d^{3}\mathbf{r}\int d^{3}\mathbf{r}^{\prime}(\mathbf{r})\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r})\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r^{\prime}})\frac{V_{\sigma\sigma}(\mathbf{r}-\mathbf{r^{\prime}})}{2}\hat{\psi}_{\sigma}(\mathbf{r^{\prime}})\hat{\psi}_{\sigma}(\mathbf{r})$
-$=$ $\sum_{\mathbf{i}_{\sigma},\mathbf{j}_{\sigma},\mathbf{k}_{\sigma},\mathbf{l}_{\sigma}}\frac{1}{2}\int d^{3}\mathbf{r}\int d^{3}\mathbf{r}^{\prime}w_{\mathbf{i}_{\sigma}}^{*}(\mathbf{r})w_{\mathbf{j}_{\sigma}}^{*}(\mathbf{r^{\prime}})\frac{4\pi\hbar^{2}a_{\sigma\sigma}}{\mu}\delta(\mathbf{r}-\mathbf{r^{\prime}})w_{\mathbf{k}_{\sigma}}(\mathbf{r^{\prime}})w_{\mathbf{l}_{\sigma}}(\mathbf{r})\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{j}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{k}_{\sigma}}\hat{b}_{\mathbf{l}_{\sigma}}$
-$=$ $\sum_{\mathbf{i}_{\sigma}}\frac{U_{\mathbf{i}_{\sigma}\mathbf{i}_{\sigma}}}{2}\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{i}_{\sigma}}\hat{b}_{\mathbf{i}_{\sigma}}+\sum_{\left\langle\mathbf{i}_{\sigma},\mathbf{j}_{\sigma}\right\rangle}\frac{U_{\mathbf{i}_{\sigma}\mathbf{j}_{\sigma}}}{2}\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{i}_{\sigma}}\hat{b}_{\mathbf{j}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{j}_{\sigma}}$ (S9)
+$H_{\mathrm{intra},\sigma}=$ $\int d^{3}\mathbf{r}\int d^{3}\mathbf{r}^{\prime}\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r})\hat{\psi}_{\sigma}^{\dagger}(\mathbf{r^{\prime}})\frac{V_{\sigma\sigma}(\mathbf{r}-\mathbf{r^{\prime}})}{2}\hat{\psi}_{\sigma}(\mathbf{r^{\prime}})\hat{\psi}_{\sigma}(\mathbf{r})$
+$=$ $\sum_{\mathbf{i}_{\sigma},\mathbf{j}_{\sigma},\mathbf{k}_{\sigma},\mathbf{l}_{\sigma}}\frac{1}{2}\int d^{3}\mathbf{r}\int d^{3}\mathbf{r}^{\prime}\left[w_{\mathbf{i}_{\sigma}}^{*}(\mathbf{r})w_{\mathbf{j}_{\sigma}}^{*}(\mathbf{r^{\prime}})\frac{4\pi\hbar^{2}a_{\sigma\sigma}}{\mu}\right.$
+$\left.\delta(\mathbf{r}-\mathbf{r^{\prime}})w_{\mathbf{k}_{\sigma}}(\mathbf{r^{\prime}})w_{\mathbf{l}_{\sigma}}(\mathbf{r})\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{j}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{k}_{\sigma}}\hat{b}_{\mathbf{l}_{\sigma}}\right]$ (8)
+$=$ $\sum_{\mathbf{i}_{\sigma}}\frac{U_{\mathbf{i}_{\sigma}\mathbf{i}_{\sigma}}}{2}\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{i}_{\sigma}}\hat{b}_{\mathbf{i}_{\sigma}}+\sum_{\left\langle\mathbf{i}_{\sigma},\mathbf{j}_{\sigma}\right\rangle}\frac{U_{\mathbf{i}_{\sigma}\mathbf{j}_{\sigma}}}{2}\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{i}_{\sigma}}\hat{b}_{\mathbf{j}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{j}_{\sigma}}$
 where
-$U_{\mathbf{i}_{\sigma}\mathbf{i}_{\sigma}}\equiv\frac{g_{\sigma\sigma}}{2}\int d^{3}\mathbf{r}\left|w_{\mathbf{i}_{\sigma}}(\mathbf{r})\right|^{4},$ (S10)
+$U_{\mathbf{i}_{\sigma}\mathbf{i}_{\sigma}}\equiv\frac{g_{\sigma\sigma}}{2}\int d^{3}\mathbf{r}\left|w_{\mathbf{i}_{\sigma}}(\mathbf{r})\right|^{4},$ (9)
 and
-$U_{\mathbf{i}_{\sigma}\mathbf{j}_{\sigma}}\equiv\frac{g_{\sigma\sigma}}{2}\int d^{3}\mathbf{r}\left|w_{\mathbf{i}_{\sigma}}(\mathbf{r})\right|^{2}\left|w_{\mathbf{j}_{\sigma}}(\mathbf{r})\right|^{2}.$ (S11)
+$U_{\mathbf{i}_{\sigma}\mathbf{j}_{\sigma}}\equiv\frac{g_{\sigma\sigma}}{2}\int d^{3}\mathbf{r}\left|w_{\mathbf{i}_{\sigma}}(\mathbf{r})\right|^{2}\left|w_{\mathbf{j}_{\sigma}}(\mathbf{r})\right|^{2}.$ (10)
 Here $g_{\sigma\sigma}=4\pi\hbar^{2}a_{\sigma\sigma}/\mu$ denotes the intra-species contact interaction strength.
 Accordingly, the interaction Hamiltonian between different components takes the form that
-$H_{\mathrm{inter}}=\sum_{\langle\mathbf{i}_{\uparrow},\mathbf{j}_{\downarrow}\rangle}U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}\hat{n}_{\mathbf{i}_{\uparrow}}\hat{n}_{\mathbf{j}_{\downarrow}},$ (S12)
+$H_{\mathrm{inter}}=\sum_{\langle\mathbf{i}_{\uparrow},\mathbf{j}_{\downarrow}\rangle}U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}\hat{n}_{\mathbf{i}_{\uparrow}}\hat{n}_{\mathbf{j}_{\downarrow}},$ (11)
 where
-$U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}\equiv\frac{g_{\uparrow\downarrow}}{2}\int d^{3}\mathbf{r}\left|w_{\mathbf{i}_{\uparrow}}(\mathbf{r})\right|^{2}\left|w_{\mathbf{j}_{\downarrow}}(\mathbf{r})\right|^{2},$ (S13)
+$U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}\equiv\frac{g_{\uparrow\downarrow}}{2}\int d^{3}\mathbf{r}\left|w_{\mathbf{i}_{\uparrow}}(\mathbf{r})\right|^{2}\left|w_{\mathbf{j}_{\downarrow}}(\mathbf{r})\right|^{2},$ (12)
 with $g_{\uparrow\downarrow}=4\pi\hbar^{2}a_{\uparrow\downarrow}/\mu$ being the inter-species contact interaction strength.
 Considering the 2D case, we approximate the Wannier function $w_{\mathbf{i}_{\sigma}}(\mathbf{r})$ as a harmonic oscillator wave function of the lowest energy level, therefore, the Wannier function (at site $\mathbf{0}$ for instance) in the 2D plane reads
-$w(x,y)=\sqrt{2\kappa}e^{-\pi\kappa(x^{2}+y^{2})},$ (S14)
+$w(x,y)=\sqrt{2\kappa}e^{-\pi\kappa(x^{2}+y^{2})},$ (13)
 where $\kappa\equiv\sqrt{\mu V_{0}/(2\hbar^{2}d^{2})}$ with $V_{0}$ being the lattice depth and $d$ being the lattice constant. As a result, the interlayer interaction strength can be expressed as
-$U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}=\frac{g_{\uparrow\downarrow}}{2}\kappa e^{-\pi\kappa\left(\Delta_{x}^{2}+\Delta_{y}^{2}\right)},$ (S15)
+$U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}=\frac{g_{\uparrow\downarrow}}{2}\kappa e^{-\pi\kappa\left(\Delta_{x}^{2}+\Delta_{y}^{2}\right)},$ (14)
 where $\Delta_{x}^{2}+\Delta_{y}^{2}\equiv|\mathbf{r_{\mathbf{i}_{\uparrow}}}-\mathbf{r_{\mathbf{j}_{\downarrow}}}|^{2}$, which depends on the twisted angle $\theta$.
 Similarly, the strength of the interactions takes the form that
-$\begin{cases}U_{\mathbf{i}_{\sigma}\mathbf{i}_{\sigma}}=&\frac{g_{\sigma\sigma}}{2}\kappa\\ U_{\mathbf{i}_{\sigma}\mathbf{j}_{\sigma}}=&\frac{g_{\sigma\sigma}}{2}\kappa e^{-\pi\kappa d^{2}}\\ U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}=&\frac{g_{\uparrow\downarrow}}{2}\kappa e^{-\pi\kappa\left(\Delta_{x}^{2}+\Delta_{y}^{2}\right)}\end{cases}.$ (S16)
+$\begin{cases}U_{\mathbf{i}_{\sigma}\mathbf{i}_{\sigma}}=&\frac{g_{\sigma\sigma}}{2}\kappa\\ U_{\mathbf{i}_{\sigma}\mathbf{j}_{\sigma}}=&\frac{g_{\sigma\sigma}}{2}\kappa e^{-\pi\kappa d^{2}}\\ U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}=&\frac{g_{\uparrow\downarrow}}{2}\kappa e^{-\pi\kappa\left(\Delta_{x}^{2}+\Delta_{y}^{2}\right)}\end{cases}.$ (15)
 In summary, the Hamiltonian of the system reads
@@
 $+\sum_{\sigma=\uparrow,\downarrow}\sum_{\langle\mathbf{i}_{\sigma},\mathbf{j}_{\sigma}\rangle}\frac{U_{\mathbf{i}_{\sigma}\mathbf{j}_{\sigma}}(g_{\sigma\sigma})}{2}\hat{n}_{\mathbf{i}_{\sigma}}\hat{n}_{\mathbf{j}_{\sigma}}$
-$+\sum_{\langle\mathbf{i}_{\uparrow},\mathbf{j}_{\downarrow}\rangle}U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}(g_{\uparrow\downarrow},\theta)\hat{n}_{\mathbf{i}_{\uparrow}}\hat{n}_{\mathbf{j}_{\downarrow}}.$ (S17)
+$+\sum_{\langle\mathbf{i}_{\uparrow},\mathbf{j}_{\downarrow}\rangle}U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}(g_{\uparrow\downarrow},\theta)\hat{n}_{\mathbf{i}_{\uparrow}}\hat{n}_{\mathbf{j}_{\downarrow}}.$ (16)
 where $\hat{n}_{\mathbf{i}_{\sigma}}\equiv\hat{b}_{\mathbf{i}_{\sigma}}^{\dagger}\hat{b}_{\mathbf{i}_{\sigma}}$ is the particle number operator that counts the number of atoms of species $\sigma$ on site $\mathbf{i}$ of the $\sigma$-layer and $U\equiv U_{\mathbf{i}_{\sigma}\mathbf{i}_{\sigma}}$.
-We investigate the physics of this system by calculating its ground state via the bosonic Gutzwiller variational approach (Krauth et al., 1992; Jaksch et al., 1998; Lanatà et al., 2012).
+We investigate the physics of this system by calculating its ground state via the bosonic Gutzwiller variational approach Krauth et al.
+(1992); Jaksch et al.
+(1998); Lanatà et al.
+(2012).
 Within the framework of mean-field theory, we use the bosonic Gutzwiller variational approach to investigate the ground state properties of the system, with the variational ground state assumed to be the site-factorized form
-$|\mathrm{GW}\rangle=|\phi_{1_{\sigma}}\rangle_{1_{\sigma}}\otimes\ldots\otimes|\phi_{N_{\mathrm{lat},\sigma}}\rangle_{N_{\mathrm{lat},\sigma}}.$ (S18)
+$|\mathrm{GW}\rangle=|\phi_{1_{\sigma}}\rangle_{1_{\sigma}}\otimes\ldots\otimes|\phi_{N_{\mathrm{lat},\sigma}}\rangle_{N_{\mathrm{lat},\sigma}}.$ (17)
 Here, $N_{\mathrm{lat},\sigma}$ is the total number of the monolayer lattice sites in $\sigma$-layer and $|\phi_{\mathbf{i}_{\sigma}}\rangle_{\mathbf{\mathbf{i}_{\sigma}}}=\sum_{n=0}^{\infty}c_{n}^{(\mathbf{\mathbf{i}_{\sigma})}}|n\rangle_{\mathbf{\mathbf{i}_{\sigma}}}$ is the local wave function at site $\mathbf{i}$ of $\sigma$-layer with $|n\rangle_{\mathbf{i}_{\sigma}}$ being the corresponding local occupation number state and $c_{n}^{(\mathbf{\mathbf{i}_{\sigma})}}$ being the variational parameter.
@@
 $=$ $\sum_{\mathbf{i}_{\sigma},\sigma=\uparrow,\downarrow}\left[\frac{U}{2}\frac{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{i}_{\sigma})}\right|^{2}n(n-1)}{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{i}_{\sigma})}\right|^{2}}+\sum_{\mathbf{j}_{\sigma}}\frac{U_{\mathbf{i}_{\sigma}\mathbf{j}_{\sigma}}}{2}\left(\frac{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{i}_{\sigma})}\right|^{2}n}{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{i}_{\sigma})}\right|^{2}}\right)\left(\frac{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{j}_{\sigma})}\right|^{2}n}{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{j}_{\sigma})}\right|^{2}}\right)\right.$
-$\left.-J\sum_{\mathbf{j}_{\sigma}}\left(\frac{\sum_{n=0}^{\infty}c_{n}^{*(\mathbf{i}_{\sigma})}c_{n+1}^{(\mathbf{i}_{\sigma})}\sqrt{n+1}}{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{i}_{\sigma})}\right|^{2}}\right)\left(\frac{\sum_{n=0}^{\infty}c_{n}^{(\mathbf{j}_{\sigma})}c_{n+1}^{*(\mathbf{j}_{\sigma})}\sqrt{n+1}}{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{j}_{\sigma})}\right|^{2}}\right)\right]$ (S19)
+$\left.-J\sum_{\mathbf{j}_{\sigma}}\left(\frac{\sum_{n=0}^{\infty}c_{n}^{*(\mathbf{i}_{\sigma})}c_{n+1}^{(\mathbf{i}_{\sigma})}\sqrt{n+1}}{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{i}_{\sigma})}\right|^{2}}\right)\left(\frac{\sum_{n=0}^{\infty}c_{n}^{(\mathbf{j}_{\sigma})}c_{n+1}^{*(\mathbf{j}_{\sigma})}\sqrt{n+1}}{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{j}_{\sigma})}\right|^{2}}\right)\right]$ (18)
 $+\sum_{\langle\mathbf{i}_{\uparrow},\mathbf{j}_{\downarrow}\rangle}U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}\left(\frac{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{i}_{\uparrow})}\right|^{2}n}{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{i}_{\uparrow})}\right|^{2}}\right)\left(\frac{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{j}_{\downarrow})}\right|^{2}n}{\sum_{n=0}^{\infty}\left|c_{n}^{(\mathbf{j}_{\downarrow})}\right|^{2}}\right).$
 During the minimization process, we have tried different initial states to obtain the global energy minimum.
-II Phase diagrams for integer filling
-Figure S1: Phase diagrams and typical configurations of the MI-MI case and SF-MI case.
+Figure 4: Phase diagrams and typical configurations of the MI-MI case and SF-MI case.
 (a) Phase diagram with respect to the interlayer interaction strength $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ and the hopping amplitude $J/U(g_{\sigma\sigma})$ at the filling with $\rho_{\uparrow}=1$ and $\rho_{\downarrow}=1$ (twist angle is fixed at $\theta=36.87^{\circ}$).
@@
 (c) and (d) Real-space particle density and superfluid order parameter distributions of the system that correspond to the red dots and the blue squares in (a) and (b), respectively.
+Appendix B Phase diagrams for integer filling
 In the main text, we have focused on the SF-SF case, where both components have non-integer fillings, rather than on the MI-MI case (both components with integer fillings) or the SF-MI case (one component with a non-integer filling and the other with an integer filling).
 This choice is due to the large excitation gap caused by the strong interactions in the integer-filling component, which hinders atom tunneling and the formation of rich phases, unlike in the SF-SF case.
-Here, we present the phase diagrams and typical configurations for the MI-MI and SF-MI cases (see Fig.
-S1).
-Similar to the conventional Bose-Hubbard model, when the interlayer interaction strength $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ is small, the components with integer filling favor the homogeneous superfluid phase and Mott-insulator phase [see Figs.
-S1 (c1) and (c2)], while the component with non-integer filling favor the homogeneous superfluid phase [see Figs.
-S1 (d1) and (d2)].
-In the MI-MI case, even when $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ is large, the density distributions of both components remain almost homogeneous due to the excitation gap [see Figs.
-S1 (c3) and (c4)].
-Conversely, in the SF-MI case, a large $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ drives the component with non-integer filling into an induced Mott-insulator phase [see Fig.
-S1 (d4)].
-Moreover, in the regime of relatively large $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ (where interlayer interactions dominate over intralayer interactions), the intense competition between the integer-filling components due to the interlayer interaction tends to make the system more prone to forming a superfluid phase.
-However, in regions dominated by interlayer interactions, stronger interlayer interactions make it easier to form a Mott-insulator phase [see Fig.
-S1 (a)].
-In the SF-MI scenario, the high-filling component can create an effective potential that localizes the low-filling component, thereby facilitating the formation of the Mott-insulator phase [see Fig.
-S1 (b)].
+Here, we present the phase diagrams and typical configurations for the MI-MI and SF-MI cases (see Fig 4).
+Similar to the conventional Bose-Hubbard model, when the interlayer interaction strength $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ is small, the components with integer filling favor the homogeneous superfluid phase and Mott-insulator phase [see Figs 4 (c1) and (c2)], while the component with non-integer filling favors the homogeneous superfluid phase [see Figs 4 (d1) and (d2)].
+In the MI-MI case, even when $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ is large, the density distributions of both components remain almost homogeneous due to the excitation gap [see Figs 4 (c3) and (c4)].
+Conversely, in the SF-MI case, a large $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ drives the component with non-integer filling into an induced Mott-insulator phase [see Fig 4 (d4)].
+Additionally, the phase boundary in Fig 4 (a) reflects the system’s evolution as the interlayer interaction increases: from being dominated by intralayer interactions, transitioning to a regime where interlayer and intralayer interactions are in balanced competition, and finally to a regime dominated by interlayer interactions.
+In the SF-MI scenario, the high-filling component can create an effective potential that localizes the low-filling component, thereby facilitating the formation of the Mott-insulator phase as $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ increases [see Fig 4 (b)].
+Figure 5: (a) Phase diagram with respect to the interlayer interaction strength $g_{\uparrow\downarrow}/g_{\sigma\sigma}$ and the hopping amplitude $J/U(g_{\sigma\sigma})$ with $\rho_{\uparrow}=3/5$ and $\rho_{\downarrow}=3/5$ (the twist angle $\theta$ is fixed at $\theta=7.63^{\circ}$ and the system size is $37\times 37$).
+(b) Real-space particle density and superfluid order parameter distributions of the system that correspond to the red dot in (a).
+(c) Similar to (a), but the system size is $49\times 49$.
+(d) Real-space particle density and superfluid order parameter distributions of the system that correspond to the red dot in (c).
+Four degenerate ground states
+[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+$E=56.56U$ $E=56.56U$ $E=56.56U$ $E=56.56U$
+Metastable States
+[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+$E=60.26U$ $E=58.20U$ $E=57.47U$ $E=56.72U$
+Table 1: Degenerate ground state and metastable state configurations for the system parameter that corresponds to Fig. 3(d2) in the manuscript.
+Appendix C Finite-size effects
+In this section, we present a detailed analysis of the finite-size effects in our numerical simulations.
+Given the spatial dependence of the interaction $U_{\mathbf{i}_{\uparrow}\mathbf{j}_{\downarrow}}(g_{\uparrow\downarrow},\theta)$ and the potential impact of the system size on the observed physical properties, we have carried out additional simulations for larger lattice sizes to ensure the reliability of our results.
+To investigate the finite-size effects, we compare the phase diagrams and typical physical quantities between two system sizes: $37\times 37$ and $49\times 49$.
+Specifically, we focus on the smallest twist angle $\theta=7.63^{\circ}$ studied in our manuscript.
+Figs 5 (a) and (c) shows a direct comparison of the phase diagrams for the $37\times 37$ and $49\times 49$ systems, and both phase diagrams display qualitatively consistent results.
+In addition to the phase diagrams, we also compare the spatial configurations of relevant physical observables, such as the superfluid order parameter $\phi_{\mathbf{i}_{\sigma}}\equiv\langle\hat{b}_{\mathbf{i}_{\sigma}}\rangle$ and density distribution $\langle\hat{n}_{\mathbf{i}_{\sigma}}\rangle$.
+Figs 5 (b) and (d) display representative configurations for both system sizes.
+The spatial profiles of the superfluid order parameter and density exhibit negligible differences between the two system sizes, further confirming that the 37×37 system adequately captures the essential physics of the system.
+Appendix D Spontaneously breaking of $C_{4}$ symmetry and metastable states
+In this section, we investigate the spontaneous breaking of $C_{4}$ rotational symmetry and the metastable states observed in our simulations.
+These phenomena arise due to the intrinsic disorder potential, leading to the four degenerate ground states and a rich set of metastable states.
+To demonstrate the degeneracy of these ground states, we performed simulations with different initial conditions for the system parameters corresponding to Fig.
+3(d2) of the main text.
+As shown in the first row of the plots in Table 1, we identified four distinct ground state configurations possessing identical energy.
+This confirms the spontaneous breaking of the $C_{4}$ symmetry in the system.
+In addition to the four degenerate ground states, our simulations reveal the presence of numerous metastable configurations within the Bose glass phase.
+These metastable states have higher energies than the ground states.
+The second row of the plots in Table 1 provides a representative set of metastable configurations for the same parameter, illustrating the diversity of particle distributions and their associated energies.
+This observation aligns with the expected behavior of a Bose glass, which is typically characterized by the coexistence of multiple metastable configurations.
+Figure 6: Identification of the Bose glass phase.
+(a) Typical configurations of the Bose glass phase [induced Mott insulator phase with non-percolated superfluid islands as well as the Fig.
+2(c4) in the main text].
+(b) Typical configurations of the induced Mott insulator phase with percolated superfluid region [Fig.
+2(c3) in the main text].
+(c) Typical configurations of the superfluid phase [Fig.
+2(c2) in the main text].
+The percolated superfluid regions across the length scale of the system size (the red square) are marked by the small red circles.
+Appendix E Bose glass identification
+To identify the Bose glass phase in our numerical simulations, we implemented a systematic approach based on the percolation properties of the superfluid regions and the spatial characteristics of the density distribution.
+This section outlines the methodology used to identify the Bose glass phase.
+We use the approach outlined by references Niederle and Rieger (2013); Johnstone et al.
+(2021b) to analyze percolation of SF clusters on a discrete function $S_{\mathbf{i}_{\sigma}}$.
+We label the lattice site $\mathbf{i}_{\sigma}$ with integer density distribution $\langle\hat{n}_{\mathbf{i}_{\sigma}}\rangle$ and zero SF order parameter $\phi_{\mathbf{i}_{\sigma}}$ as MI site with $S_{\mathbf{i}_{\sigma}}=0$, and label the site with non-integer filling and nonzero $\phi_{\mathbf{i}_{\sigma}}$ as SF site with $S_{\mathbf{i}_{\sigma}}=1$.
+The BG phase is identified by examining whether a percolated superfluid region exists on the length scale of the system size.
+If none of the connected clusters formed by the SF sites ($S_{\mathbf{i}_{\sigma}}=1$) percolates across the system, the phase is identified as the BG phase.
+This phase is characterized by isolated SF clusters embedded within an MI sea ($S_{\mathbf{i}_{\sigma}}=0$), resulting in a non-vanishing density of SF sites that lack global connectivity Niederle and Rieger (2013).
+Indeed, in the typical BG phase observed in our numerical simulations, isolated superfluid islands (i.e., non-percolated superfluid regions) are usually surrounded by MI regions.
+The “induced MI” phase defined in the main text is characterized by regions of the system where sites are filled with an integer number of atoms and have no associated SF order parameter.
+Consequently, the BG phase can still be considered a form of “induced MI” as long as there are MI regions present.
+Following the percolation identification approach on a square lattice described in references Niederle and Rieger (2013); Barman et al.
+(2013), we apply a similar method in our system with an external cylinder box trap.
+Specifically, we inscribe the largest possible square within the circular region to assess percolation.
+To ensure robustness, this inscribed square is systematically rotated from $0{{}^{\circ}}$ to $360{{}^{\circ}}$, allowing us to evaluate percolation for all orientations.
+Examples of the identification of BG phase are presented Fig.
+6.